### v1

# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

Abstract.
We consider a Dirichlet series $D_{F,G}(s)$ attached to two modular forms $F$ and $G$ of an orthogonal group of real signature $(2,4)$, involving their Fourier–Jacobi coefficients. When $F$ is a Hecke eigenform and $G$ a Poincaré series, our main result gives that $D_{F,G}(s)$ is equal to the standard $L$-function attached to $F$, up to an explicit constant. To establish this, we use a correspondence between binary Hermitian forms and ideals of quaternion algebras, as established by Latimer, together with the fact that the even Clifford algebra of a three-dimensional definite quadratic space can be identified with a quaternion division algebra. Our work should be seen as a generalisation of a work of Kohnen and Skoruppa, whose result corresponds to the case of the orthogonal group of real signature $(2,3)$.
Acknowledgements.
Part of this work was completed during the second-named author’s PhD studies. This author was supported by the HIMR/UKRI "Additional Programme for Mathematical Sciences", grant number EP/V521917/1, as well as by a scholarship from the Onassis Foundation.
2020 Mathematics Subject Classification:
11F55, 11E88, 11E41, 11F66, 11M41
1. Introduction
Let $F$ and $G$ be two degree two Siegel cusp forms of even weight $k\geq 0$. In their groundbreaking work [KS89], Kohnen and Skoruppa define and study the Dirichlet series
$D_{F,G}(s):=\zeta(2s-2k+4)\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
where $\phi_{N}$ and $\psi_{N}$ are the $N$th Fourier–Jacobi coefficients of $F$ and $G$ respectively, and $\langle\,$$\rangle$ is the Petersson inner product on Jacobi forms of weight $k$ and index $N$. The significance of their work derives mainly from the fact that if one further assumes that $F$ is a Hecke eigenform and $G$ is in the Maass space, then this Dirichlet series is proportional to $Z_{F}(s)$, the spinor $L$-function attached to $F$. Actually, as explained in their paper, it is sufficient to establish such a relation for the case where $G$ is taken to be the lift of a Poincaré series, as such lifts generate the Maass space in the symplectic setting. In particular, they show in [KS89, Theorem 3],
$D_{F,\mathcal{P}_{k,D}}(s)=A(Q)Z_{F}(s),$
where $Q$ denotes any quadratic form of discriminant $D$ representing $1$, and $A(Q)$ is the $Q$th Fourier coefficient of $F$. Here, $\mathcal{P}_{k,D}$ is a lift of a classical Poincare series associated to $k$ and $D$ (see [KS89, page 552]).
In this paper, we obtain a result analogous to Kohnen and Skoruppa’s for modular forms of orthogonal groups of signature (2,4). We will first describe our result and then explain how it fits into a more general picture, alongside Kohnen and Skoruppa’s work.
Let $S$ denote an even, positive definite symmetric matrix of rank $2$, such that $\mathbb{Z}^{2}$ is a maximal lattice with respect to $S$. We put $K:=\mathbb{Q}\left(\sqrt{-\det S}\right)$. Let
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\,\,\,S_{1}=\begin{pmatrix}&&1\\ &-S_{0}&\\ 1&&\end{pmatrix}$,
and write $(V_{i},\phi_{i})$ for the quadratic spaces associated to $S_{i}$, for $i=0,1$. We also write $\textup{SO}^{\phi_{i}}(V_{i})$ for the corresponding special orthogonal groups. Let now $F$ denote an orthogonal cusp form of integer even weight $k\geq 0$ with respect to a congruence subgroup $\Gamma_{S}$ of $\textup{SO}^{\phi_{1}}(V_{1})$. Assume $\{\phi_{m}\}_{m=1}^{\infty}$ are its Fourier–Jacobi coefficients. We also consider a Maass lift of a Poincaré series, given by
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{k,D,r})(\tau,z)e(N\tau^{\prime}),$
where $V_{N}$ is an index-raising operator acting on Fourier–Jacobi forms of lattice index and $(D,r)\in\mathbb{Z}\times\mathbb{Z}^{2}$ is in the support of the maximal lattice $\mathbb{Z}^{2}$ with respect to $S$. Here, $(\tau^{\prime},z,\tau)$ is in the associated symmetric space (see 5). For any finite set of primes $\mathcal{P}$, we then consider the Dirichlet series
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\,\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s},$
and set
$D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s),$
where $\langle\ ,\ \rangle$ denotes the inner product on the space of Fourier–Jacobi forms of weight $k$ and (lattice) index $NS$, and $\zeta_{K}$ denotes the Dedekind zeta function attached to $K$.
We now take $D=-\ell q$, with $\ell\in\mathbb{Z}_{>0}$ and $q$ the level of $S$. We let $L_{0}:=\mathbb{Z}^{4}$ and then choose $\xi=(1,0,0,\ell)\in V_{0}$, for which $\phi_{0}[\xi]=\ell$ and $2\phi_{0}(\xi,L_{0})=\mathbb{Z}$. The main Theorem of this paper is the following.
Theorem 1.1.
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. Assume $F$ is a Hecke eigenform for the corresponding Hecke algebra. Then, outside a finite set of primes $\mathcal{P}$, we have
$D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=A(\xi)L_{\mathcal{P}}(F;s-k+2),$
where $L(F;s)$ is the standard $L$-function attached to $F$, and $A(\xi)$ is the $\xi^{th}$ Fourier coefficient of $F$. In particular, if $A(\xi)\neq 0$, then $D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)$ is proportional to $L_{\mathcal{P}}(F;s-k+2)$.
We mention here that we also have computational evidence which suggests that Proposition 7.1 can be extended to also include the case of $K=\mathbb{Q}(\sqrt{-2})$. However, we have not yet been able to prove it, and we hope to return to this in the near future.
Let us now explain how the above results can be seen in a more general setting and how this paper is related to the recent work [Psy25] of one of the authors. It is a well-known fact that there is an accidental isogeny of the symplectic group $\textup{Sp}_{2}(\mathbb{R})$ and of $\textup{SO}(2,3)$, the special orthogonal group of signature $(2,3)$. Moreover, the spinor $L$-function attached to $F$ above can be identified with the standard $L$-function attached to a holomorphic modular form on an appropriate orthogonal group of real signature $(2,3)$, as for example is discussed in [Shi04, Section 25]. In fact, there are more accidental isogenies between classical and orthogonal groups. Examples are $\textup{U}(2,2)$ and $\textup{SO}(2,4)$, and $\textup{Sp}(2,\mathbb{H})$ and $\textup{SO}(2,6)$, where $\mathbb{H}$ denotes the skew field of real quaternions. Modular forms can be defined for orthogonal groups of any real signature $(2,n+2)$, $n\geq 1$. Crucially, the associated symmetric space is of tube type and has the structure of a Shimura variety. It is therefore natural to ask whether the method of Kohnen and Skoruppa can be generalised to the orthogonal setting. This has been the subject of study in a recent work of one of the authors in [Psy25].
Indeed, consider the same setting as above, but with $S$ now having rank $n\geq 1$. We denote by $H(\xi)$ the subgroup of $\textup{SO}^{\phi_{0}}(V_{0})$ given by
$H(\xi)_{\mathbb{Q}}:=\{g\in\textup{SO}^{\phi_{0}}(V_{0})\mid g\xi=\xi\},$
which is a negative definite orthogonal group of rank $n+1$. It is one of the main results by Shimura in [Shi04] that the congruence subgroup of $\textup{SO}^{\phi_{0}}(V_{0})$ which fixes the lattice $L_{0}:=\mathbb{Z}^{n+2}$ acts on the set
$\left\{x\in V_{0}\mid\phi_{0}(x,x)=-D/q\textup{ and }2\phi_{0}(x,L_{0})=\mathbb{Z}\right\},$
and there are finitely many equivalence classes under this action. Let $\{\xi_{i}\}_{i=1}^{h}$ be representatives for this action. These representatives correspond to elements of the finite set
$H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C),$ (1)
where $C=\{x\in\textup{SO}^{\phi_{0}}(V_{0})_{\mathbb{A}}\mid xL_{0}=L_{0}\}$. Assume $\{u_{i}\}_{i=1}^{h}$ are these elements. This is the analogue of the main theorem for binary quadratic forms of fixed discriminant.
Let also $\{f_{j}\}_{j=1}^{h}$ denote an orthonormal basis of simultaneous eigenforms on the set (1). Denote by $L(-,s)$ the standard $L$-function attached to either $F$ or any of the $f_{j}^{\prime}s$. We then have the following Theorem.
Theorem 1.2.
([Psy25, Theorem 8.2]) Outside a finite set of primes $\mathcal{P}$, $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ can be written as
$L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+(n+3)/2\right)^{-1}\times\\ \times\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i})\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases}.$
Here, $\displaystyle{\zeta_{\xi_{i}}(s)=\sum_{d=1}^{\infty}n(\xi_{i};d)d^{-s}}$, with
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\Gamma(L_{0})\right\}.$
and the $A_{f_{j}}$ are expressions depending on $f_{j}$ and the Fourier coefficients of $F$. For any zeta function, the subscript $\mathcal{P}$ means that we do not take into account the terms sharing factors with elements of $\mathcal{P}$.
It is therefore clear that to obtain a clear-cut Euler product, it is necessary to relate
$\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i}),$
to the standard $L$-function attached to $\overline{f_{j}}$ for each $j=1,\cdots,h$. In [Psy25, Section 9], we restricted ourselves to orthogonal groups for which $h=1$ and obtained clear-cut Euler products in [Psy25, Theorem 9.2, Theorem 9.3].
The main contribution of the present paper is to drop the condition $h=1$ for the $n=2$ case. Our strategy is to exploit the isogeny between the unitary group $\textup{U}(2,2)$ and $\textup{SO}(2,4)$. In particular, we identify the $\xi_{i}$’s with classes of binary Hermitian forms of discriminant $\ell$. According to Latimer in [Lat35], these correspond to specific classes of ideals in a certain quaternion algebra $B$, which allow us to identify each $\zeta_{\xi_{i}}(s)$ as the partial zeta function associated with classes of ideals in $B$. We then identify such a quaternion algebra as an even Clifford algebra associated to $\textup{SO}(3)$, and interpret the $f_{j}$’s as lifts to modular forms on the corresponding Clifford group. Our result then follows using Shimura’s results in [Shi04, Section 22.11].
Finally, we should mention here that Gritsenko in [Gri92] has considered the analogous problem for the case of Hermitian modular forms of degree two over the imaginary quadratic field $K=\mathbb{Q}(i)$. Gritsenko’s approach is to use factorisation methods in parabolic Hecke rings, and, in fact, is quite different from the techniques used in this paper. Indeed, our work is closer to the approach of Kohnen and Skoruppa [KS89], where the representation numbers of binary Hermitian forms in our work play an analogue role to the representation numbers of binary quadratic forms in their work. In fact, we also establish a relation of the standard $L$-function of $\textup{SO}(3)$ with the zeta functions $\zeta_{\xi_{i}}$’s (Theorem 10.5). These zeta functions count solutions of congruences and hence can potentially be used to compute the $L$-function explicitly. We do believe that our main result is compatible with Grisenko’s main Theorem in [Gri92, (5.3)], but we have not yet checked all the details. We hope to return to this in the near future.
Notation.
We denote the space of $m\times n$ matrices with coefficients in a ring $R$ with $\textup{M}_{m,n}(R)$. If $n=m$, we often use the notation $\textup{M}_{n}(R)$. By $1_{n}$, we denote the $n\times n$ identity matrix. For any matrix $M\in\textup{M}_{n}(R)$, we denote by $\det(M),\textup{ tr}(M)$ the determinant and trace of $M$ respectively. By $\textup{GL}_{n}(R)$, we denote the matrices in $M_{n}(R)$ with non-zero determinant and by $\textup{SL}_{n}(R)$ the matrices with determinant $1$. For any vector $v$, we denote by $v^{t}$ its transpose. We also use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$. By $\textup{diag}(A_{1},A_{2},\cdots,A_{n})$, we will denote the block diagonal matrix with the matrices $A_{1},A_{2},\cdots,A_{n}$ in the diagonal blocks. For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. Finally, let $\zeta(s)$ denote the Riemann zeta function.
2. Preliminaries
Let $F$ denote the field $\mathbb{Q}$ or $\mathbb{Q}_{p}$ for any rational prime $p$. Let $V$ denote a finite-dimensional vector space over $F$ and let $\phi:V\times V\longrightarrow F$ be a non-degenerate symmetric bilinear form. We call $(V,\phi)$ a quadratic space. We put $\phi[x]:=\phi(x,x)$ for all $x\in V$. We denote by $A(V,\phi)$ or simply by $A(V)$ the Clifford algebra associated to $(V,\phi)$. We define the canonical automorphism $x\longmapsto x^{\prime}$ and the canonical involution $x\longmapsto x^{*}$ of $A(V)$ by the property $v=-v^{\prime}=v^{*}$ for all $v\in V$. We then set
$A^{+}(V):=\{\alpha\in A(V)\mid\alpha^{\prime}=\alpha\},$
$A^{-}(V):=\{\alpha\in A(V)\mid\alpha^{\prime}=-\alpha\}.$
We also define the special orthogonal and Clifford group by
$\textup{SO}^{\phi}(V):=\{\alpha\in\textup{GL}(V)\mid\phi[\alpha x]=\phi[x],\ \forall x\in V\},$
$G(V):=\{\alpha\in A(V)^{\times}\mid\alpha V\alpha^{-1}=V\},$
respectively, and put $G^{+}(V):=G(V)\cap A^{+}(V)$, the even Clifford group. We also define the homomorphisms
$\displaystyle\begin{split}\tau:G^{+}(V)&\longrightarrow\textup{SO}^{\phi}(V)\\ \alpha&\longmapsto\tau(\alpha)x:=\alpha x\alpha^{-1}.\end{split}$ (2)
and
$\displaystyle\begin{split}\nu:G^{+}(V)&\longrightarrow F^{\times}\\ \alpha&\longmapsto\alpha\alpha^{*}.\end{split}$ (3)
It is well known that $\tau$ is surjective with kernel $F^{\times}$.
Let now $\mathfrak{g}$ denote a maximal order in $F$, i.e. $\mathfrak{g}=\mathbb{Z}$ or $\mathbb{Z}_{p}$, depending on $F=\mathbb{Q}$ or $\mathbb{Q}_{p}$. By a $\mathfrak{g}$-lattice $L$ in $V$, we mean a finitely generated $\mathfrak{g}$-module that spans $V$ over $F$. It is called $\mathfrak{g}$-integral if $\phi[x]\in\mathfrak{g}$ for all $x\in L$. We say $L$ is $\mathfrak{g}$-maximal if it maximal among all integral $\mathfrak{g}$-lattices. We also define its dual lattice via
$L^{*}:=\{x\in V\mid 2\phi(x,y)\in\mathfrak{g},\ \forall y\in L\}.$
The level $q$ of $L$ is defined as the least positive integer such that $q\phi[x]\in\mathfrak{g},\ \forall x\in L^{*}$. We also define the integral orthogonal group via
$\textup{SO}^{\phi}(L):=\{\alpha\in\textup{SO}^{\phi}(V)\mid\alpha L=L\}.$
By $A(L)$ we denote the subring of $A(V)$ which is generated by $L$ and $\mathfrak{g}$. We also put $A^{+}(L):=A(L)\cap A^{+}(V)$.
For any $q\in\mathbb{Q}^{\times}$ and $\mathfrak{b}$ a fractional ideal of $\mathbb{Q}$, we define
$L[q,\mathfrak{b}]:=\{x\in V\mid\phi[x]=q\textup{ and }\phi(x,L)=\mathfrak{b}\}.$ (4)
Finally, given two lattices $L,M$ in $V$, we denote by $[L/M]$ the fractional ideal generated by $\det\alpha$ for all $\alpha\in\textup{GL}(V)$ such that $\alpha L\subset M$.
3. Modular Forms on $\textup{SO}(2,n)$
Let now $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$ (both identified with the standard basis). Assume $S$ is an even integral positive definite symmetric matrix of rank $n$. Here, even means $S[x]:=x^{t}Sx\in 2\mathbb{Z}$ for all $x\in L$. We define
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\ S_{1}:=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
of real signatures $(1,n+1)$ and $(2,n+2)$, respectively. Let also $V_{0}:=\mathbb{Q}^{n+2}$ and $V_{1}:=\mathbb{Q}^{n+4}$ and consider the quadratic spaces $(V_{0},\phi_{0}),\ (V_{1},\phi_{1})$, where
$\displaystyle\phi_{i}:V_{i}\times V_{i}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{i}y,$
for $i=0,1$. We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just $(x,y)\longmapsto-x^{t}Sy/2$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}$-lattice with respect to $\phi$.
From [Shi04, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}$-maximal lattice in $V_{0}$.
If now $K\supset\mathbb{Q}$ is a field, the corresponding orthogonal groups of $K$-rational points are given by
$\textup{SO}^{\phi_{0}}(V_{0}\otimes K):=\{g\in\textup{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},$
$\textup{SO}^{\phi_{1}}(V_{1}\otimes K):=\{g\in\textup{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
Let now $\mathcal{H}_{S}$ denote one of the connected components of
$\{Z\in V_{0}\otimes_{\mathbb{Q}}\mathbb{C}\mid\phi_{0}[\textup{Im}Z]>0\}.$
In particular, if we let
$\mathcal{P}_{S}:=\{y^{\prime}=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\,\,\phi_{0}[y^{\prime}]>0\},$
we choose
$\mathcal{H}_{S}:=\{z=x+iy\in V_{0}\otimes_{\mathbb{R}}\mathbb{C}\mid y\in\mathcal{P}_{S}\}.$ (5)
For a matrix $g\in M_{n+4}(\mathbb{R})$, we write it as
$g=\begin{pmatrix}\alpha&a^{t}&\beta\\ b&A&c\\ \gamma&d^{t}&\delta\end{pmatrix}$,
with $A\in M_{n+2}(\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$ (6)
gives a well-defined transitive action of $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})^{+}$ on $\mathcal{H_{S}}$, where $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})^{+}$ denotes the connected component of the identity of $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})$. The denominator of the above expression is the factor of automorphy
$j(g,Z):=-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta.$
Let now $L_{1}:=\mathbb{Z}^{n+4}$ and define the groups
$\Gamma(L_{0}):=\{g\in\textup{SO}^{\phi_{0}}(V_{0})\mid gL_{0}=L_{0}\},$
$\Gamma(L_{1}):=\{g\in\textup{SO}^{\phi_{1}}(V_{1})\mid gL_{1}=L_{1}\}.$
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap\textup{SO}^{\phi_{0}}(V_{0}\otimes\mathbb{R})^{+}$. Moreover, we set
$\Gamma_{S}:=G_{\mathbb{R}}^{0}\cap\Gamma(L_{1}),$
and
$\widehat{\Gamma}_{S}:=\{M\in\Gamma_{S}\mid M\in 1_{n+4}+\textup{M}_{n+4}(\mathbb{Z})S_{1}\},$
the discriminant kernel.
Definition 3.1.
Let $k\in\mathbb{Z}$ and $\widehat{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ a subgroup of finite index. A holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k$ with respect to $\Gamma$ if it satisfies the equation
$(F|_{k}\gamma)(Z):=j(\gamma,Z)^{-k}F(\gamma\langle Z\rangle)=F(Z)$
for all $\gamma\in\Gamma$ and $Z\in\mathcal{H}_{S}$. We will denote the set of such forms by $M_{k}(\Gamma)$.
Now, if $\widehat{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$, $F\in M_{k}(\Gamma)$ admits a Fourier expansion of the form (see [Sug95, (5.10)])
$F(Z)=\sum_{r\in L_{0}^{*}}A(r)e(r^{t}S_{0}Z).$ (7)
It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ ($\overline{\mathcal{P}_{S}}$ denotes the closure of $\mathcal{P_{S}}$, see [Sch22, Theorem 1.5.2]). From [Sch22, Theorem 1.6.23], and because $L$ is maximal, we have the following definition for cusp forms.
Definition 3.2.
If $\widehat{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$, $F\in M_{k}(\Gamma)$ is called a cusp form if it admits a Fourier expansion of the form
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z),$
We denote the space of cusp forms by $S_{k}(\Gamma)$.
Let now $S_{k}(L,\phi)$ denote the vector space of Fourier–Jacobi cusp forms of weight $k$ and (lattice) index $L$ (see [Moc19, Definition 1.23, 1.26]).
If $F\in S_{k}(\Gamma)$, with $\widehat{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ and we write $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$ with $\tau^{\prime},\tau\in\mathbb{C},\ z\in\mathbb{C}^{n}$, we have that for any $m\in\mathbb{Z}$,
$F(\tau^{\prime}+m,z,\tau)=F(\tau^{\prime},z,\tau).$
Hence, we can write
$F(Z)=\sum_{m\geq 1}\phi_{m}(\tau,z)e^{2\pi im\tau^{\prime}},$ (8)
and we call the functions $\phi_{m}(\tau,z)$ the Fourier–Jacobi coefficients of $F$. We note that then $\phi_{m}\in S_{k}(L,m\phi)$ (see [Sug95, (5.10), (5.11)]).
We will now define a special class of Fourier–Jacobi forms, called Poincaré series. To begin with, we define the support of the lattice $L=\mathbb{Z}^{n}$ to be
$\textup{supp}(L,\phi):=\left\{(D,r)\mid D\in\mathbb{Q}_{\leq 0},r\in L^{*},D\equiv\frac{1}{2}S[r]\pmod{\mathbb{Z}}\right\}$.
Now, if we write $r\longmapsto S^{-1}r$ with $r\in L$, we get
$qD\equiv\frac{1}{2}qS^{-1}[r]\pmod{q\mathbb{Z}}$,
where $q$ is the level of the lattice. This then implies $D\in\frac{1}{q}\mathbb{Z}$. Hence, by writing $D\longmapsto D/q$, we get that equivalently
$\textup{supp}(L,\phi)=\left\{(D/q,S^{-1}r)\mid D\in\mathbb{Z}_{\leq 0},r\in L,D\equiv\frac{1}{2}qS^{-1}[r]\pmod{q}\right\}$.
Let then
$\widetilde{\textup{supp}}(L,\phi):=\{(D,r)\in\mathbb{Z}_{\leq 0}\times L\mid(D/q,S^{-1}r)\in\textup{supp}(L,\sigma)\}.$ (9)
and in the following this is the set we will use.
If now $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$, we define the Poincaré series $P_{k,D,r}$ as in [Moc19, Definition 2.2]. If $k>n/2+2$, $P_{k,D,r}$ is absolutely and uniformly convergent on compact subsets of $\mathbb{H}\times(L\otimes\mathbb{C})$ ($\mathbb{H}$ the usual upper half plane) and defines an element of $S_{k}(L,\phi)$ (see [Moc19, Theorem 2.3, (i)]).
Finally, if $N\geq 1$, we denote by $V_{N}:S_{k}(L,\phi)\longrightarrow S_{k}(L,N\phi)$ the index-raising operator defined in [Sug95, Lemma 6.1] and [Moc19, Definition 4.25]. In particular, for $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$, we define
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{k,D,r})(\tau,z)e(N\tau^{\prime}).$ (10)
From [Sug95, Corollary 6.7], we have that $\mathcal{P}_{k,D,r}\in S_{k}(\widehat{\Gamma}_{S})$ and this is actually an element of the Maass space, as is shown by Sugano in [Sug95, Corollary 6.7]. Here, by abusing notation, we write $P_{k,D,r}$ to actually denote the Poincaré series $P_{k,D,r}/\lambda_{k,D}$, where $\lambda_{k,D}$ is defined in [Moc19, Theorem 2.3, (i)].
4. Dirichlet series
We will now focus on the $n=2$ case. Assume $k>3$ even and $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$. We take $F\in S_{k}(\Gamma_{S})$ and consider the Dirichlet series
$\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s):=\sum_{N=1}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s},$ (11)
where $\langle\ ,\ \rangle$ denotes the inner product on the space of Fourier–Jacobi cusp forms of weight $k$ and lattice index (see [Moc19, Definition 1.33]). As in [Psy25, Lemma 6.1], this converges absolutely and uniformly on compact subsets of $\mathbb{C}$ for $\textup{Re}(s)>k+1$.
We can now rewrite $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ as in [Psy25, Proposition 7.3], but with a small difference. Let us make this explicit. Let
$\widehat{\Gamma}(L_{0}):=\{M\in\Gamma(L_{0})\mid M\in 1_{n+2}+\textup{M}_{n+2}(\mathbb{Z})S_{0}\},$ (12)
and write $\widehat{\Gamma}^{+}(L_{0})$ for the connected component of the identity inside $\widehat{\Gamma}(L_{0})$.
From now on, we choose $D=-\ell q$, with $\ell$ a positive integer and $q$ the level of $S$.
Lemma 4.1.
We have that $\widehat{\Gamma}(L_{0})$ acts on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ (see equation (4)) and there is a finite number of orbits under this action.
Proof.
We defer the proof of this to the next Section. ∎
Let now $h_{1}:=\#\left(L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})\right)$. In a similar fashion to the discussion before [Psy25, Proposition 7.3], we fix a set of representatives $\{\xi_{i}\}_{i=1}^{h_{1}}$ for the orbits of this action, such that $\xi_{i}\in\mathcal{P}_{S}$ for all $i$. We also define
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}.$ (13)
We then have the following Proposition, as in [Psy25, Proposition 7.3].
Proposition 4.2.
Let $(-\ell q,r)\in\widetilde{\textup{supp}}(L,\phi)$ and $\mathcal{P}$ be any finite set of primes, which includes the prime factors of $D$. We put
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s},$
which converges absolutely for $\textup{Re}(s)>k+1$ by comparison with $D_{F,\mathcal{P}_{k,D,r}}(s)$. For each $i=1,\cdots,h_{1}$, let also
$\zeta_{\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi_{i};N)N^{-s}.$ (14)
We then have that
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)D_{F,\xi_{i},\mathcal{P}}(s),$ (15)
where $D_{F,\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}$, and $h_{1}=\#\left(L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})\right)$.
5. Equivalence of Quadratic Spaces
Our aim is now to understand the Dirichlet series $D_{F,\xi_{i},\mathcal{P}}(s)$. To that end, in this Section, we will introduce an isomorphism of the quadratic space $(V_{0},\phi_{0})$ with a quadratic space of Hermitian matrices over an imaginary quadratic field. Let
$S:=\begin{cases}\begin{pmatrix}2&0\\ 0&2m\end{pmatrix}\textup{ if }m\equiv 1,2\pmod{4}\\ \begin{pmatrix}2&1\\ 1&(m+1)/2\end{pmatrix}\textup{ if }m\equiv 3\pmod{4}\end{cases}$. (16)
for a square-free integer $m$, and set $K:=\mathbb{Q}\left(\sqrt{-\textup{det}S}\right)$. We note that when $K$ is of class number one (this is the case for which we will give our main result), by the theory of quadratic forms, there is only one class of quadratic forms with $\mathbb{Z}^{2}$ as a maximal lattice. That is, the choice of $S$ is canonical in the sense that any other choice would give us an isomorphic quadratic space, with the same arithmetic structure.
Let us now write $\mathcal{O}_{K}$ for the ring of integers of $K$. We have $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, where
$\omega=\begin{cases}\sqrt{-m}\,\,\,\,\textup{ if }m\equiv 1,2\pmod{4}\\ \frac{1+\sqrt{-m}}{2}\,\,\,\,\textup{ if }m\equiv 3\pmod{4}\end{cases}$. (17)
We now consider the four dimensional (over $\mathbb{Q}$) quadratic space $(U,\psi)$, where
$U:=\left\{h\in M_{2}(K)\mid\overline{h}^{t}=h\right\},$
and $\psi[h]=\det(h),\ \forall h\in U$. One has (see for example [Shi06b, (2.1d)]) that the corresponding bilinear form is given by $2\psi(x,y)=\textup{tr}(xy^{\iota})$, where $\iota$ here is the main involution on $M_{2}(K)$. We define the following elements of $U$
$e:=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\,\,f:=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},\,\ g:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$,
and set $\Lambda:=\mathbb{Z}e+\mathbb{Z}f+\mathcal{O}_{K}g$. By [Shi06b, p. 495] we have that this is a $\mathbb{Z}$-maximal lattice in $U$.
Let also $U^{+}:=\{u\in U\mid u>0\}$, the set of positive definite matrices in $U$. We then have the following Lemma.
Lemma 5.1.
There is an isomorphism of quadratic spaces $f_{\omega}:(U,\psi)\longrightarrow(V_{0},\phi_{0})$, given as
[…]
where $x,y,z,w\in\mathbb{Q}$. Moreover, the map $f_{\omega}$ is a bijection between $U^{+}$ and $\mathcal{P}_{S}$.
Proof.
One sees easily that the map $f_{\omega}$ is well defined linear isomorphism of the $\mathbb{Q}$-vector spaces $U$ and $V_{0}$ (compare also with [Wer19, Lemma 1.24 and Lemma 1.25]). Moreover, an easy computation gives $\phi_{0}[f_{\omega}(u)]=\psi[u]$, for $u\in U$, and therefore $f_{\omega}$ is indeed an isometry. The last statement of the Lemma follows from Sylvester’s criterion. ∎
It is now clear that $f_{\omega}(\Lambda)=L_{0}$. Moreover, $f_{\omega}$ induces the following isomorphism of orthogonal groups:
$\displaystyle\begin{split}g_{\omega}:\textup{SO}^{\psi}(U)&\xrightarrow[]{\sim}\textup{SO}^{\phi_{0}}(V_{0})\\ \alpha&\longmapsto f_{\omega}\circ\alpha\circ f_{\omega}^{-1}\end{split}$ (19)
This also extends to an isomorphism of the integral groups $\textup{SO}^{\psi}(\Lambda)\xrightarrow[]{\sim}\textup{SO}^{\phi_{0}}(L_{0})$. Indeed, if $\alpha\in\textup{SO}^{\psi}(\Lambda)$, we have
$g_{\omega}(\alpha)(L_{0})=f_{\omega}\circ\alpha\circ f_{\omega}^{-1}(L_{0})=f_{\omega}\circ\alpha(\Lambda)=f_{\omega}(\Lambda)=L_{0},$
using the fact that $\alpha(\Lambda)=\Lambda$ and that $f_{\omega}$ is an isomorphism.
We now have the following Lemma, for the elements of $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$, where $\ell$ is a non-zero integer.
Lemma 5.2.
With the notation as above, we have
$h\in\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]\iff f_{\omega}(h)\in L_{0}\left[\ell,2^{-1}\mathbb{Z}\right].$
Proof.
From Lemma 5.1, we have $\psi[h]=\ell\iff\phi_{0}[f_{\omega}(h)]=\ell$. Therefore, we need to check only the condition on the ideals. Let
$h=\begin{pmatrix}a&b+\omega c\\ b+\overline{\omega}c&d\end{pmatrix}\in U$
with $a,b,c,d\in\mathbb{Q}$. From [Shi06b, (2.18)], we have that the condition on the ideals is equivalent to
$a\mathbb{Z}+d\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}((b+\overline{\omega}c)\mathcal{O}_{K})=\mathbb{Z}.$ (20)
We will now consider two cases:
Case 1: $m\equiv 3\pmod{4}$. Then
$\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-(2b+c)\mathbb{Z}-\left(b+\frac{m+1}{2}c\right)\mathbb{Z}=\mathbb{Z}.$ (21)
But $\mathcal{O}_{K}$ has basis $\{1,\omega\}$. In order to compute $\textup{Tr}_{K/\mathbb{Q}}((b+\overline{\omega}c)\mathcal{O}_{K})$, we will compute the traces of the linear maps $x\longmapsto\alpha x$ for $\alpha=b+\overline{\omega}c$ and $\alpha=(b+\overline{\omega}c)\omega=\frac{m+1}{4}c+b\omega$.
$\textup{Tr}_{K/\mathbb{Q}}(b+\overline{\omega}c)=\textup{Tr}\left(\begin{pmatrix}b+c&(m+1)c/4\\ -c&b\end{pmatrix}\right)=2b+c.$
$\textup{Tr}_{K/\mathbb{Q}}\left((m+1)c/4+b\omega\right)=\textup{Tr}\left(\begin{pmatrix}(m+1)c/4&-(m+1)b/4\\ b&b+(m+1)c/4\end{pmatrix}\right)=b+\frac{m+1}{2}c.$
We now see that (20) and (21) are equivalent.
Case 2: $m\equiv 1,2\pmod{4}$. In this case,
$\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-2b\mathbb{Z}-2mc\mathbb{Z}=\mathbb{Z}.$ (22)
In a similar fashion to before, we compute
$\textup{Tr}_{K/\mathbb{Q}}(b+\overline{\omega}c)=\textup{Tr}\left(\begin{pmatrix}b&mc\\ -c&b\end{pmatrix}\right)=2b.$
$\textup{Tr}_{K/\mathbb{Q}}((b+\overline{\omega}c)\omega)=\textup{Tr}\left(\begin{pmatrix}mc&-mb\\ b&mc\end{pmatrix}\right)=2mc.$
Again, it is now evident that (20) and (22) are equivalent. ∎
The next step is to see how the action of $\widehat{\Gamma}(L_{0})$ on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ translates in the Hermitian setting. We will first recall some basic definitions related to unitary groups of degree 2. We first define the Hermitian upper half plane
$\mathbb{H}_{2}:=\{Z\in M_{2}(\mathbb{C})\mid-i(Z-\overline{Z}^{t})>0\}.$
We fix an embedding $K\xhookrightarrow{}\mathbb{C}$. The special unitary group of degree 2 over $K$ is given by
$\textup{SU}(2,2;K):=\{g\in\textup{SL}_{4}(K)\mid\overline{g}^{t}Jg=J\},$
where $J:=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Now, $M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\textup{SU}(2,2;K)$ acts on $\mathbb{H}_{2}$ via
$Z\longmapsto M\langle Z\rangle:=(AZ+B)(CZ+D)^{-1}.$
Finally, we set $\textup{SU}(2,2;\mathcal{O}_{K}):=\textup{SU}(2,2;K)\cap M_{4}(\mathcal{O}_{K})$. We then have the following Proposition.
Proposition 5.3.
There is an isomorphism $\Psi:\textup{SL}_{2}(\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}^{+}(L_{0})$ such that
$f_{\omega}(\gamma u\gamma^{*})=\Psi(\gamma)f_{\omega}(u),\ \forall\gamma\in\textup{SL}_{2}(\mathcal{O}_{K}),\ u\in U_{+},$ (23)
where $\gamma^{*}:=\overline{\gamma}^{t}$ and $U^{+}$ is the set of positive definite matrices in $U$.
Proof.
From [Wer19, Theorem 3.17], we have that there is an isomorphism $\Phi:\textup{SU}(2,2;\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}_{S}$. Moreover, there is an embedding $\epsilon:\textup{SL}_{2}(\mathcal{O}_{K})\xhookrightarrow{}\textup{SU}(2,2;\mathcal{O}_{K})$ by sending $\gamma\longrightarrow\textup{diag}(\gamma,\overline{\gamma}^{-t})$, which is compatible with the embedding $\eta:\widehat{\Gamma}^{+}(L_{0})\xhookrightarrow{}\widehat{\Gamma}_{S}$ via $\gamma\longmapsto\textup{diag}(1,\gamma,1)$, because of the explicit form of $\Phi$ given in [Wer19, Theorem 3.13]. We define a map $\Psi:\textup{SL}_{2}(\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}^{+}(L_{0})$ by setting $\Psi(\gamma)$ the element such that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, and we claim this is an isomorphism with the required property. First of all, it is a homomorphism from the way it is defined. To check $\Psi$ is injective, assume $\Psi(\gamma_{1})=\Psi(\gamma_{2})$. Then $\Phi(\epsilon(\gamma_{1}))=\Phi(\epsilon(\gamma_{2}))$ and therefore $\epsilon(\gamma_{1})=\pm\epsilon(\gamma_{2})$, which in turn implies $\gamma_{1}=\pm\gamma_{2}$, as wanted. Now, if $M\in\widehat{\Gamma}^{+}(L_{0})$, we have $\eta(M)\in\widehat{\Gamma}_{S}$ and so there exists $\widetilde{\gamma}\in\textup{SU}(2,2;\mathcal{O}_{K})$ such that $\Phi(\widetilde{\gamma})=\eta(M)$. Now, let
$\widetilde{\gamma}=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$,
with $A,B,C,D\in M_{2}(\mathcal{O}_{K})$. From [Wer19, Theorem 3.13], we must have $\det A=\det D=1$ and $f_{\omega}(A^{\sharp}B)=f_{\omega}(AC^{\sharp})=f_{\omega}(BD^{\sharp})=f_{\omega}(C^{\sharp}D)=0$, where denotes the adjugate of a matrix (see [Wer19, (3.9)]). But $f_{\omega}$ is injective and so $B=C=0$ and then $\widetilde{\gamma}=\epsilon(A)$, with $A\in\textup{SL}_{2}(\mathcal{O}_{K})$. Therefore, $\Psi$ is also surjective and hence an isomorphism.
Let now $u\in U^{+}$. We then have that $iu\in\mathbb{H}_{2}$. From [Wer19, Theorem 3.5], we have $f_{\omega}(M\langle Z\rangle)=\Phi(M)\langle f_{\omega}(Z)\rangle$ for all $Z\in\mathbb{H}_{2}$ and $M\in\textup{SU}(2,2;\mathcal{O}_{K})$. By taking now $M=\epsilon(\gamma)$ with $\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})$ and using that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, (23) follows by unwinding the definitions of the actions.∎
We end this Section with the following identification of the orthogonal and Clifford groups of $(U,\psi)$, as is given by Shimura in [Shi06b, Proposition 2.2].
Proposition 5.4.
Let $\mathcal{G}:=\{\alpha\in\textup{GL}_{2}(K)\mid\det(\alpha)\in\mathbb{Q}^{\times}\}$. There is an isomorphism $\Xi$ of $G^{+}(U)$ onto $\mathcal{G}$, such that if $\alpha\in G^{+}(U)$ and $\xi=\Xi(\alpha)$, then $\nu(\alpha)=\det(\xi)$ and $\tau(\alpha)h=\det(\xi)^{-1}\xi h\overline{\xi}^{t}$ for all $h\in U$. In particular, $\textup{SO}^{\psi}(U)$ is isomorphic to $\mathcal{G}/\mathbb{Q}^{\times}$.
6. The Dirichlet Series $D_{F,\xi_{i},\mathcal{P}}(s)$
In this Section, we will relate $D_{F,\xi_{i},\mathcal{P}}(s)$ of Proposition 4.2 to $L$-functions. We will first establish Lemma 4.1, namely that $\widehat{\Gamma}(L_{0})$ acts on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ and there is a finite number of orbits under this action. For $\ell\in\mathbb{Z}_{>0}$, we define
$L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]:=\{x=(x_{1},x_{2},x_{3},x_{4})\in L_{0}[\ell,2^{-1}\mathbb{Z}]\mid x_{1}>0\}.$
$L_{0}^{-}[\ell,2^{-1}\mathbb{Z}]:=\{x=(x_{1},x_{2},x_{3},x_{4})\in L_{0}[\ell,2^{-1}\mathbb{Z}]\mid x_{1}<0\}.$
Note that $\phi_{0}[x]=\ell\implies 2x_{1}x_{4}-S[x]=2\ell\implies x_{1}x_{4}>0$, so $x_{4}$ has the same sign as $x_{1}$.
We claim that there is an equivalence between $L_{0}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}(L_{0})$ and $L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}^{+}(L_{0})$, where we recall that $\widehat{\Gamma}^{+}(L_{0})$ is the connected component of the identity inside $\widehat{\Gamma}(L_{0})$. First of all
$L_{0}[\ell,2^{-1}\mathbb{Z}]=L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]\sqcup L_{0}^{-}[\ell,2^{-1}\mathbb{Z}].$
If $\sigma:=\textup{diag}(-1,1,1,-1)\in\widehat{\Gamma}^{+}(L_{0})\backslash\widehat{\Gamma}(L_{0})$, we have that $\widehat{\Gamma}(L_{0})=\widehat{\Gamma}^{+}(L_{0})\sqcup\widehat{\Gamma}^{+}(L_{0})\sigma$. Hence, the element $\sigma$ interchanges the two sets $L_{0}^{\pm}[\ell,2^{-1}\mathbb{Z}]$, and therefore, the claim follows.
Moreover, using Lemma 5.2, Proposition 5.3, and the fact that the action of $\textup{SL}_{2}(\mathcal{O}_{K})$ on $\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]$ factors through $\textup{SL}_{2}(\mathcal{O}_{K})/\{\pm 1\}$, we have further the identification of $L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}^{+}(L_{0})$ and $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$, where $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]:=\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]\cap U^{+}$.
The quotient $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ is studied in detail by Shimura in [Shi06b]. In particular, we consider $\xi:=\textup{diag}(1,\ell)\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$. We define
$W:=\{u\in U\,\,|\,\,\psi(u,\xi)=0\},$ (24)
and put $H(\xi):=\textup{SO}^{\psi}(W)$, where we understand that $\psi$ is restricted to $W$.
We will now introduce localised and adelised groups and lattices, as usual. The reader can refer to [Shi97, Chapter 8] to recall the necessary definitions.
We have the inclusion of quadratic spaces $(W,\psi)\subset(U,\psi)$, hence the embeddings $A^{+}(W)\subset A^{+}(U)$, and $G^{+}(W)\subset G^{+}(U)$. For each rational prime $p$, let
$C_{p}:=\{g\in\textup{SO}^{\psi}(U)_{p}\,\,|\,\,g\Lambda_{p}=\Lambda_{p}\},$ (25)
$J_{p}:=\{\gamma\in G^{+}(U)_{p}\,\,|\,\,\tau(\gamma)\in C_{p},\,\,\nu(\gamma)\in\mathbb{Z}_{p}^{\times}\},$ (26)
and put $J:=G^{+}(U)_{\infty}\prod_{p}J_{p}$ and $C:=\textup{SO}^{\psi}(U)_{\infty}\prod_{p}C_{p}$.
Shimura’s work in [Shi06b] gives the following Proposition.
Proposition 6.1.
There is a bijection
[…]
In particular, the set on the right-hand side is finite.
Proof.
First of all, there is an equivalence of the sets $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ and $\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]/\Delta$, where $\Delta:=\{\alpha\in\textup{GL}_{2}(\mathcal{O}_{K})\mid\det(\alpha)=\pm 1\}$, as is shown in [Shi06b, p. 500]. Moreover, from the proof of [Shi06b, Theorem 2.10], the set $\Lambda[\ell,2^{-1}\mathbb{Z}]/\Delta$ is in bijection with $G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$, because the class number of $\mathbb{Q}$ is $1$. ∎
Therefore, Lemma 4.1 follows from the above discussion and this Proposition.
In order to now understand the Dirichlet series $D_{F,\xi_{i},\mathcal{P}}(s)$, we will work back in the quadratic space $(V_{0},\phi_{0})$. Note that $f_{\omega}$ of Lemma 5.1 induces an isomorphism between the Clifford algebras $A(U)$ and $A(V_{0})$, which then gives an isomorphism of the Clifford groups $G^{+}(U)$ and $G^{+}(V_{0})$. Moreover, $g_{\omega}$ of (19) is naturally extended to an isomorphism between $\textup{SO}^{\psi}(U)_{\mathbb{A}}$ and $\textup{SO}^{\phi_{0}}(V_{0})_{\mathbb{A}}$. We set
$\widetilde{W}:=f_{\omega}(W)=\{x\in V_{0}\mid\phi_{0}(x,f_{\omega}(\xi))=0\},$
[…]
$\widetilde{C}_{p}:=g_{\omega}(C_{p})=\{g\in\textup{SO}^{\phi_{0}}(V_{0})_{p}\mid gL_{0,p}=L_{0,p}\},$
$\widetilde{J}_{p}:=f_{\omega}(J_{p})=\{\gamma\in G^{+}(V_{0})_{p}\mid\tau(\gamma)\in\widetilde{C}_{p},\ \nu(\gamma)\in\mathbb{Z}_{p}^{\times}\},$
and put $\widetilde{J}:=G^{+}(V_{0})_{\infty}\prod_{p}\widetilde{J}_{p}$ and $\widetilde{C}:=\textup{SO}^{\phi_{0}}(V_{0})_{\infty}\prod_{p}\widetilde{C}_{p}$.
We now have the following Proposition.
Proposition 6.2.
We work in the quadratic space $(V_{0},\phi_{0})$. We have the following commutative diagram, where the vertical arrows are bijections.
[…]
Here, $\iota$ maps $\widehat{\Gamma}(L_{0})v\longmapsto\Gamma(L_{0})v$ for all $v\in L_{0}[\ell,2^{-1}\mathbb{Z}]$ and $\kappa,\lambda$ are the bijections defined by Shimura in [Shi06b, Theorem 1.6, (ii)] and [Shi06a, Theorem 2.2, (ii)], respectively.
Proof.
Let us first recall how $\kappa$ and $\lambda$ are defined.
From [Shi06b, Theorem 1.6, (ii)], given $\epsilon\in G^{+}(\widetilde{W})_{\mathbb{A}}$, we take $\alpha\in G^{+}(V_{0})$ such that $\epsilon\in\alpha\widetilde{J}$ and set $\kappa(\epsilon):=\widehat{\Gamma}(L_{0})\tau(\alpha)^{-1}\xi$.
Similarly, from [Shi06a, Theorem 2.2, (ii)], given $\epsilon^{\prime}\in\widetilde{H}(\xi)_{\mathbb{A}}$, we take $\alpha^{\prime}\in\textup{SO}^{\phi_{0}}(V_{0})$ such that $\epsilon^{\prime}\in\alpha^{\prime}\widetilde{C}$ and set $\lambda(\epsilon^{\prime}):=\Gamma(L_{0})(\alpha^{\prime})^{-1}\xi$.
Let now $\epsilon\in G^{+}(\widetilde{W})_{\mathbb{A}}$. On the one hand, we have $\lambda(\tau(\epsilon))=\Gamma(L_{0})\alpha^{-1}\xi$, where $\alpha\in\textup{SO}^{\phi_{0}}(V_{0})$ such that $\tau(\epsilon)\in\alpha\widetilde{C}$. On the other hand, $\iota(\kappa(\epsilon))=\iota\left(\widehat{\Gamma}(L_{0})\tau(\beta)^{-1}\xi\right)=\Gamma(L_{0})\tau(\beta)^{-1}\xi$, where $\beta\in G^{+}(V_{0})$ such that $\epsilon\in\beta\widetilde{J}$.
We now write $\tau(\epsilon)=\alpha c$, with $c\in\widetilde{C}$ and $\epsilon=\beta j$, with $j\in\widetilde{J}$. Then $\tau(\epsilon)=\tau(\beta)\tau(j)$ and so $\tau(\beta)^{-1}\alpha=\tau(j)c^{-1}\in\widetilde{C}$. Hence $\tau(\beta)^{-1}\alpha\in\textup{SO}^{\phi_{0}}(V_{0})\cap\widetilde{C}=\Gamma(L_{0})$, i.e. $\tau(\beta)^{-1}\in\Gamma(L_{0})\alpha^{-1}$. It then follows that $\lambda(\tau(\epsilon))=\iota(\kappa(\epsilon))$, as wanted. ∎
We are now ready to deal with $D_{F,\xi_{i},\mathcal{P}}(s)$. This is done in [Psy25, Section 8] but we need to modify it slightly.
We observe that for each $i=1,\cdots,h_{1}$, the series $D_{F,\xi_{i},\mathcal{P}}(s)$ depends only on the $\Gamma^{+}(L_{0})$-equivalence class of $\xi_{i}$ (because of the invariance of $F$ under $\Gamma_{S}$, see also [Sch22, p. 26]). Since $\widehat{\Gamma}^{+}(L_{0})\leq\Gamma^{+}(L_{0})$, we may think the map $\iota$ of Proposition 6.2 as
$\iota:L_{0}^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}^{+}(L_{0})\longrightarrow L_{0}^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\Gamma^{+}(L_{0}).$
Therefore, we can write (15) as
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)D_{F,\iota(\xi_{i}),\mathcal{P}}(s).$ (27)
[…]
$M(u_{i};\xi)_{p}:=\widetilde{H}(\xi)_{p}\cap u_{i}\widetilde{C}u_{i}^{-1}\textup{ and }M(u_{i};\xi)_{f}:=\prod_{p}M(u_{i};\xi)_{p}.$
Define then $e(\xi)_{i}:=\#\{\widetilde{H}(\xi)_{\mathbb{Q}}\cap M(u_{i};\xi)_{f}\}$ for $1\leq i\leq h_{2}$, and $\displaystyle{\mu(\xi):=\sum_{i=1}^{h_{2}}e(\xi)^{-1}_{i}}$. Furthermore, for any $f\in V(\xi)$, we set
$A_{f}:=\mu(\xi)^{-1}\sum_{i=1}^{h_{2}}\frac{\overline{f}(u_{i})}{e(\xi)_{i}}A(\xi_{i}),$ (28)
[…]
$D_{F,\iota(\xi_{i}),\mathcal{P}}(s)=\zeta_{\mathcal{P}}\left(2s-2k+4\right)^{-1}L_{\mathcal{P}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\lambda^{-1}(\iota(\xi_{i})))L_{\mathcal{P}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}},$ (29)
using the notation of Proposition 6.2. This is [Psy25, (16)] tailored to our situation.
[…]
$D_{F,\iota(\xi_{i}),\mathcal{P}}(s)=\zeta_{\mathcal{P}}\left(2s-2k+4\right)^{-1}L_{\mathcal{P}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\tau(\zeta_{i}))L_{\mathcal{P}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}}.$ (30)
7. Quaternion Algebras
Now that we have an understanding of the Dirichlet series $D_{F,\iota(\xi_{i}),\mathcal{P}}(s)$, our next step is to understand the zeta functions $\zeta_{\xi_{i}}(s)$ of (14). Our goal is to relate them to a partial zeta function in a quaternion algebra. To that end, we need several preparations. The quadratic space $(W,\psi)$ is three-dimensional and, as a result, its even Clifford algebra can be identified with a quaternion algebra, which in our case is a division algebra. Before making this explicit, let us recall some basic facts on quaternion algebras. We mainly follow [Voi21].
Let $B$ be a quaternion algebra over $\mathbb{Q}$ with a standard involution $a\longmapsto\overline{a}$. A $\mathbb{Z}$-order $\mathcal{O}\subset B$ is a $\mathbb{Z}$-lattice that is also a subring of $\mathbb{Z}$. For any $\mathbb{Z}$-lattice $I\subset B$, we define
$\mathcal{O}_{L}(I):=\{\alpha\in B\mid\alpha I\subseteq I\},\,\,\mathcal{O}_{R}(I):=\{\alpha\in B\mid I\alpha\subseteq\alpha\}.$
These are orders in $B$ ([Voi21, Lemma 10.2.7]), and are called the left and right order of $I$, respectively. We say $I$ is right-invertible if there exists lattice $I^{\prime}\subset B$ such that $II^{\prime}=\mathcal{O}_{L}(I)$. Similarly, we define the notion of a left-invertible lattice, and we say $I$ is invertible if there is a two-sided inverse $I^{\prime}\subset B$, so that
$II^{\prime}=\mathcal{O}_{L}(I)=\mathcal{O}_{R}(I^{\prime}),\,\,I^{\prime}I=\mathcal{O}_{L}(I^{\prime})=\mathcal{O}_{R}(I).$
Now, if $\alpha\in B$, we define its reduced norm as $\textup{nrd}(\alpha):=\alpha\overline{\alpha}$. We define the reduced norm $\textup{nrd}(I)$ of a $\mathbb{Z}$-lattice $I$, as the $\mathbb{Z}$-submodule of $\mathbb{Q}$ generated by $\{\textup{nrd}(\alpha)\mid\alpha\in I\}$. This is fractional ideal of $\mathbb{Q}$ ([Voi21, Lemma 16.3.2]).
Given now a $\mathbb{Z}$-order $\mathcal{O}\subset B$, we say that $I$ is a left-fractional $\mathcal{O}$-ideal if $\mathcal{O}\subseteq\mathcal{O}_{L}(I)$; similarly on the right.
Two $\mathbb{Z}$-lattices $I$ and $J$ are said to be in the same right class and write $I\sim_{R}J$, if $\exists\alpha\in B^{\times}$ such that $\alpha I=J$; similarly for the left. The relation $\sim_{R}$ defines an equivalence relation on the set of $\mathbb{Z}$-lattices in $B$ and we denote the equivalence class of a $\mathbb{Z}$-lattice $I$ by $[I]_{R}$. The (right) class set of $\mathcal{O}$ is
$\textup{Cls}_{R}(\mathcal{O}):=\left\{[I]_{R}:I\textup{ an invertible right fractional-}\mathcal{O}\textup{ ideal}\right\}.$
We say two orders $\mathcal{O},\mathcal{O^{\prime}}$ are of the same type, if $\exists\alpha\in B^{\times}$ such that $\mathcal{O}^{\prime}=\alpha^{-1}\mathcal{O}\alpha$. Note that this is true if and only if $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are isomorphic as $\mathbb{Z}$-algebras. Moreover, we say that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are locally isomorphic, if the localisations $\mathcal{O}_{p}$ and $\mathcal{O}^{\prime}_{p}$ are of the same type for all rational primes $p$. Here, we define the localisation of a $\mathbb{Z}$-lattice in $B$ as usual (see, for example, [Voi21, Section 9.4]).
We define the genus of a $\mathbb{Z}$-order $\mathcal{O}\subset B$ to be the set of all $\mathbb{Z}$-orders in $B$ which are locally isomorphic to $\mathcal{O}$. The type set of $\mathcal{O}$ is the set of isomorphism classes of orders in the genus of $\mathcal{O}$. If
$X:=B_{\infty}^{\times}\prod_{p}\mathcal{O}_{p}^{\times},$
[…]
Let us now define the quaternion algebra $B:=K+\epsilon K$ with $\epsilon^{2}=$-$\ell$ and $\epsilon a=\overline{a}\epsilon$ for all $a\in K$. Let also $\mathfrak{o}:=\mathcal{O}_{K}+\epsilon\mathcal{O}_{K}$ denote an order in $B$. From the proof of [Shi06b, Theorem 2.10], we have $\Xi(A^{+}(W))=B$ and $\Xi(G^{+}(W))=B^{\times}$, where $\Xi$ is the map of Proposition 5.4. Moreover, from the same Theorem, $\Xi(G^{+}(W)_{p}\cap J_{p})=\mathfrak{o}_{p}^{\times}$ for all $p$.
[…]
Proposition 7.1.
Let $D:=\{x\in\widetilde{H}(\xi)_{\mathbb{A}}\mid x(L_{0}\cap\widetilde{W})=L_{0}\cap\widetilde{W}\}$. Then, $\widetilde{H}(\xi)_{\mathbb{A}}\cap\widetilde{C}=D$, at least in the following cases:
  • $m\equiv 3\pmod{4}$, $\ell$ is square-free, and $\textup{gcd}(\ell,m)=1$.
  • $m\equiv 1\pmod{4}$, $\ell$ is square-free, $\textup{gcd}(\ell,m)=1$, and $\ell\equiv 1\pmod{4}$.
Proof.
Because the lattice $L_{0}$ is taken maximal in $(V_{0},\phi_{0})$, we have from [Shi06, Paragraph 6.3] that $[L_{0}^{*}/L_{0}]=\det(S_{0})\mathbb{Z}$. Moreover, as in [Psy25, Lemma 9.1], $\widetilde{H}(\xi)$ is represented via the matrix
$T:=\frac{1}{2}\begin{pmatrix}-2\ell\\ &-S\end{pmatrix}$,
with respect to a basis of $M:=\mathbb{Z}^{3}$. We will now show that under our assumptions on $k$ and $m$, $M$ is a maximal $\mathbb{Z}$–lattice.
If $m\equiv 3\pmod{4}$, then $\det(2T)=2\ell\det(S)=2\ell m$ and because of the conditions on $\ell$ and $m$, $\det(2T)/2$ is square-free. Hence, from [Sch22, Proposition 1.6.12], the lattice $\mathbb{Z}^{3}$ is maximal with respect to $(\widetilde{W},\phi_{0})$.
If $m\equiv 1\pmod{4}$, we will use [Sch22, Lemma 1.6.5 (ii)]. Assume we pick $x=(x_{1},x_{2},x_{3})\in M^{*}$ such that $T[x]\in\mathbb{Z}$. We will show that $x\in M$, which will give us that $M$ is $\mathbb{Z}$–maximal. We write $x_{1}=y_{1}/(2\ell),\ x_{2}=y_{2}/2,\ x_{3}=y_{3}/(2m)$, where $y_{i}\in\mathbb{Z}$. Then, the condition $T[x]\in\mathbb{Z}$ is equivalent to $4\ell m\mid my_{1}^{2}+\ell my_{2}^{2}+\ell y_{3}^{2}$. Now, by the assumptions, $\ell$ and $4m$ are coprime. Hence, we can look at each divisibility separately.
  • $\ell\mid my_{1}^{2}+\ell my_{2}^{2}+\ell y_{3}^{2}$. This implies $\ell\mid my_{1}^{2}$, hence $\ell\mid y_{1}$, because $\gcd(\ell,m)=1$ and $\ell$ square–free.
  • $4m\mid my_{1}^{2}+\ell my_{2}^{2}+\ell y_{3}^{2}$. First, this implies $m\mid\ell y_{3}^{2}$, so $m\mid y_{3}$, because $\gcd(\ell,m)=1$ and $m$ square–free. By now writing $y_{3}=mz_{3}$, the divisibility is equivalent to $4\mid y_{1}^{2}+\ell y_{2}^{2}+\ell mz_{3}^{2}$. But since $\ell,m\equiv 1\pmod{4}$, we have $4\mid y_{1}^{2}+y_{2}^{2}+z_{3}^{2}$, which implies all of them must be even.
These conditions imply that $x_{i}\in\mathbb{Z}$ for all $i=1,2,3$, as wanted.
Now, again from [Shi06, Paragraph 6.3], we have $[M^{*}/M]=\det(2T)\mathbb{Z}=2\ell\det(S)\mathbb{Z}$. Since $\phi_{0}[f_{\omega}(\xi)]=\ell$ and $2\phi_{0}(f_{\omega}(\xi),L_{0})=\mathbb{Z}$, we have $\phi_{0}[f_{\omega}(\xi)]\phi_{0}(f_{\omega}(\xi),L_{0})^{-2}=4\ell\mathbb{Z}$ and this equals $2[M^{*}/M][L_{0}^{*}/L_{0}]^{-1}$. Hence, from [Mur18, Corollary 4.3], we obtain $L_{0}\cap\widetilde{W}$ is $\mathbb{Z}$-maximal in $\widetilde{W}$.
This now means that $(L_{0}\cap\widetilde{W})_{p}$ is maximal in $\widetilde{W}_{p}$ for all primes $p$. Hence, from [Shi04, Proposition 11.12, (iv)], we obtain $\widetilde{H}(\xi)_{p}\cap\widetilde{C}=D_{p}$ for all primes $p$, as $t_{p}$ is always even in our case. Here, $D_{p}:=D\cap\widetilde{H}(\xi)_{p}$. Hence, similarly to [Psy25, Proposition 9.3], we obtain $D=\widetilde{H}(\xi)_{\mathbb{A}}\cap\widetilde{C}$, as required. ∎
Remark 7.1.
We believe that the case $m\equiv 2\pmod{4}$ can also be included, for appropriate choices of $\ell$. We have obtained computational evidence which suggests this, and we hope to relax this restriction in the near future.
We now have a surjective map (see [Voi21, Lemma 17.4.13])
$\displaystyle\rho:\textup{Cls}\mathfrak{o}$ $\displaystyle\longrightarrow\textup{Typ}\mathfrak{o}$
$\displaystyle[I]$ $\displaystyle\longmapsto[\mathcal{O}_{L}(I)].$
Because of the identifications we mentioned at the beginning of the Section, we can also think of this map as $\rho:B^{\times}\backslash B^{\times}_{\mathbb{A}}/X\longrightarrow B^{\times}\backslash B^{\times}_{\mathbb{A}}/N_{B}(X)$. We then have the following Proposition.
Proposition 7.2.
Assume $m$ and $\ell$ are as in Proposition 7.1. With the notation as above, we have the following commutative diagram, where the horizontal arrows are bijections:
${G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)}$${B^{\times}\backslash B^{\times}_{\mathbb{A}}/X}$${H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)}$${B^{\times}\backslash B^{\times}_{\mathbb{A}}/N_{B}(X)}$$\scriptstyle{\tau}$$\scriptstyle{\Xi}$$\scriptstyle{\rho}$$\scriptstyle{A^{+}}$
Here, for any lattice $N$, $A^{+}(N)$ is given in as in Section 2.
Proof.
[…]
On the other hand, the map $\tau$ will send the class of $a\in G^{+}(W)_{\mathbb{A}}$ to the class of the lattice $[\tau(a)(L\cap W)]$ in $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$. Then, via the map $N\longmapsto A^{+}(N)$, we have that
$A^{+}(\tau(a)(L\cap W))=A^{+}(a(L\cap W)a^{-1})=aA^{+}(L\cap W)a^{-1}=a\mathfrak{o}a^{-1},$
because $\mathfrak{o}=A^{+}(L\cap W)$. This shows that the diagram is indeed commutative. ∎
8. Representation Numbers of Binary Hermitian Forms
The aim of this Section is to relate the expressions $n(\xi_{i};d)$ of equation (13) with $\xi_{i}\in L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}(L_{0})$ to representation numbers of binary Hermitian quadratic forms. We recall,
$n(\xi_{i};d)=\#\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\},$
where $D=-\ell q$, with $\ell$ a positive integer and $q$ the level of $S$. From Lemma 5.2 and Proposition 5.3, we have seen the identification
$\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]/\textup{SL}_{2}(\mathcal{O}_{K})\longleftrightarrow L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]/\widehat{\Gamma}^{+}(L_{0})$
induced by the map $f_{\omega}$. We set $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$, for $i=1,\ldots,h_{1}$. We first prove the following Lemma.
Lemma 8.1.
Let $\Lambda=M_{2}(\mathcal{O}_{K})\cap U$, as above. We have the following two statements:
  • If $\ell\in\mathbb{Z}$, then $h\in\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]\implies h\in\Lambda$.
  • If $\ell\in\mathbb{Z}$ square-free and $\gcd(\ell,2m)=1$, then
    $\{h\in\Lambda\mid\det(h)=\ell\}=\Lambda[\ell,\mathbb{Z}]\cup\Lambda\left[\ell,2^{-1}\mathbb{Z}\right].$
Proof.
The first statement follows from [Shi04, Lemma 6.2, (3)], and the second statement is [Shi06b, Lemma 4.5]. ∎
By this Lemma, we have in particular that $f_{i}\in\Lambda$ whenever $\ell\in\mathbb{Z}$. For a $d\in\mathbb{N}$, $f\in\Lambda$ with $\textup{det}(f)=\ell$, we set
$r(f;d)=\#\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f\right\},$ (31)
where the notation $X\sim Y$ for two $2\times 2$ matrices $X$ and $Y$ means that there is a matrix $A\in\textup{SL}_{2}(\mathcal{O}_{K})$ such that $AX\overline{A}^{t}=Y$. We now prove the following Proposition.
Proposition 8.2.
Let $\ell\in\mathbb{Z}_{>0}$ and $\xi_{i}\in L_{0}^{+}[\ell,2^{-1}\mathbb{Z}]$. Recall $f_{i}:=f_{\omega}^{-1}(\xi_{i})\in\Lambda^{+}[\ell,2^{-1}\mathbb{Z}]$. Then, for any $d\in\mathbb{N}$, we have $n(\xi_{i};d)=r(f_{i};d)$.
Proof.
We consider two cases:
Case 1: $m\equiv 1,2\pmod{4}$. In this case, from (16), $S=\textup{diag}(2,2m)$, and so we have $q=4m$ and $D=-4\ell m$. Therefore, if $s=(s_{1},s_{2})$, the congruence $D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}$ can be written as
$4m\ell\equiv ms_{1}^{2}+s_{2}^{2}\pmod{4md}$,
with $s_{1}\pmod{2d},\,\,s_{2}\pmod{2md}$. Now, $m\mid s_{2}^{2}\implies m\mid s_{2}$ because $m$ is square-free. If we write $s_{2}\longmapsto ms_{2}$, the congruence becomes
$s_{1}^{2}+ms_{2}^{2}\equiv-4\ell\pmod{4d}$,
with $s_{1},s_{2}\pmod{2d}$. But since $m\equiv 1,2\pmod{4}$, $-m$ is not a quadratic residue $\pmod{4}$, so $s_{1}s_{2}$ must be both even. Hence, the congruence is equivalent to
$s_{1}^{2}+ms_{2}^{2}\equiv-\ell\pmod{d}$,
with $s_{1},\ s_{2}\pmod{d}$. Hence, by taking $h:=s_{1}+\omega s_{2}$, we see that this congruence is equivalent to $h\overline{h}\equiv-\ell\pmod{d}$ and $h\in\mathcal{O}_{K}/d\mathcal{O}_{K}$. Therefore, the map
$(s_{1},s_{2})\longmapsto h:=\frac{s_{1}}{2}+\omega\frac{s_{2}}{2m},$
induces a bijection of
$\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}$
and the set
$\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\}.$
Hence $n(\xi_{i};d)=r(f_{i};d)$.
Case 2: $m\equiv 3\pmod{4}$. In this case, $q=m$ and hence $D=-\ell m$. We use the method used in [Psy25, Section 9.3]. In particular, if
$P=\begin{pmatrix}0&1\\ -1&2\end{pmatrix},\,\,\,Q=\begin{pmatrix}1&-(m+1)/2\\ 0&1\end{pmatrix}$,
we have $PSQ=\textup{diag}(1,m)$. Now, we set $t=Ps$ and the congruence $D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}$ is equivalent to
$\frac{1}{2}qS^{-1}[P^{-1}t]\equiv D\pmod{qd}$,
where $t=(t_{1},t_{2})$ with $t_{1}\pmod{d}$ and $t_{2}\pmod{md}$. By substituting and performing the operations, we get
$mt_{1}^{2}-mt_{1}t_{2}+\frac{m+1}{4}t_{2}^{2}\equiv-\ell m\pmod{md}$,
with $t_{1},t_{2}$ as before. But $\gcd(m,(m+1)/4)=1$, so $m\mid t_{2}^{2}\implies m\mid t_{2}$, because $m$ is square-free. Hence, the congruence is equivalent to
$t_{1}^{2}-mt_{1}t_{2}+\frac{m(m+1)}{4}t_{2}\equiv-\ell\pmod{d}$,
with $t_{1},t_{2}\pmod{d}$. Hence, by taking $\displaystyle{h:=\left(t_{1}-\frac{m+1}{2}t_{2}\right)+\omega t_{2}}$ (recall that $\omega=(1+\sqrt{-m})/2$ in this case), we see that this is equivalent to $h\overline{h}\equiv-\ell\pmod{d}$ and $h\in\mathcal{O}_{K}/d\mathcal{O}_{K}$. Therefore, by changing back the variables, the map
${(s_{1},s_{2})\longmapsto h:=\left(\frac{m+1}{2}\frac{s_{1}}{m}-\frac{s_{2}}{m}\right)+\omega\frac{-s_{1}+2s_{2}}{m}}$
induces a bijection of
$\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}$
and the set
$\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\}.$
Hence $n(\xi_{i};d)=r(f_{i};d)$. ∎
Our next step is to relate the numbers $r(f_{i};d)$ to some representation numbers attached to the binary Hermitian form associated with the Hermitian matrix $f_{i}$.
Let $I_{\ell}$ denote the number of $\textup{SL}_{2}(\mathcal{O}_{K})$-equivalence classes in $\Lambda$ for the set of Hermitian binary forms of fixed discriminant $\ell$. We have $I_{\ell}\geq h_{1}$, and we extend the $\textup{SL}_{2}(\mathcal{O}_{K})$-equivalence classes $\{f_{i}\}_{i=1}^{h_{1}}$, with $f_{i}\in\Lambda[\ell,2^{-1}\mathbb{Z}]$, to a complete set of $\textup{SL}_{2}(\mathcal{O}_{K})$-equivalence classes $\{f_{i}\}_{i=1}^{I_{\ell}}$ (note that by the first statement of Lemma 8.1, we have $\Lambda[\ell,2^{-1}\mathbb{Z}]\subset\Lambda$ and hence our original $f_{i}$’s belong in $\Lambda$). Here, we will abuse the notation and write $f_{i}$ for both the binary Hermitian form and the Hermitian matrix representing the form. That is, for $u,v\in K$, we write $f_{i}(u,v):=(u,v)f_{i}(\bar{u},\bar{v})^{t}$.
For each $i=1,\cdots,I_{\ell}$, we introduce the set
$P(f_{i},d):=\{(u,v)\in\mathcal{O}_{K}\times\mathcal{O}_{K}\,\,|\,\,f_{i}(u,v)=d,\,\,\,<u,v>=\mathcal{O}_{K}\},$
where the notation $<u,v>$ is the ideal generated by $u$ and $v$ in $\mathcal{O}_{K}$. We moreover define the primitive representation numbers $p(f_{i},d):=\#P(f_{i},d)$. We define the set
$\mathcal{R}(d,-\ell):=\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\right\}$,
and following [EGM97, Definition 3.4], we define a map
$\varphi_{f_{i}}:P(f_{i},d)\rightarrow\mathcal{R}(d,-\ell),\,\,\,\varphi_{f_{i}}\left((u,v)\right):=h,$
where for any $(u,v)\in P(f_{i},d)$, we pick $r,s\in\mathcal{O}_{K}$ such that $\begin{pmatrix}r&s\\ u&v\end{pmatrix}\in\textup{SL}_{2}(\mathcal{O}_{K})$ and set $h$ to be the entry
$\begin{pmatrix}r&s\\ u&v\end{pmatrix}f_{i}\begin{pmatrix}\bar{r}&\bar{u}\\ \bar{s}&\bar{v}\end{pmatrix}=\begin{pmatrix}*&h\\ \bar{h}&d\end{pmatrix}$.
It is shown in [EGM97, Definition 3.4] that this map is well-defined and from [EGM97, Proposition 3.5] induces a map
$\varphi_{f_{i}}:P(f_{i},d)/E(f_{i})\rightarrow\mathcal{R}(d,-\ell),$
where $E(f_{i}):=\left\{\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})\,\,|\,\,\gamma f_{i}\gamma^{*}=f_{i}\right\}$ acting linearly on the set $P(f_{i},d)$. Let now
$\mathcal{R}(f_{i},d):=\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\},$
We then have $\textup{Im}(\varphi_{f_{i}})\subseteq\mathcal{R}(f_{i},d)$ and $r(f_{i},d)=\#\mathcal{R}(f_{i},d)$, where $r(f_{i},d)$ is defined in (31). Our aim is to show that $\varphi_{f_{i}}$ induces a bijection between $P(f_{i},d)/E(f_{i})$ and $\mathcal{R}(f_{i},d)$.
To that end, we form the disjoint union of the domains $P(f_{i},d)/E(f_{i})$ for all $i=1,\cdots,I_{\ell}$, and construct the map
$\varphi_{\ell}:\bigsqcup_{i\in I_{\ell}}P(f_{i},d)/E(f_{i})\longrightarrow\mathcal{R}(d,-\ell)$
such that $\varphi_{\ell}{\mid(P(f_{i},d)/E(f_{i}))}:=\varphi_{f_{i}}$ for all $i=1,\cdots,I_{\ell}$. From [EGM97, Theorem 3.7], the map $\varphi_{\ell}$ is bijective.
In particular, this shows that each $\phi_{f_{i}}$ is injective and surjective onto $\mathcal{R}(f_{i},d)$, hence proving our claim. Hence, if we set $e(f_{i}):=\#E(f_{i})$, we have that $\frac{p(f_{i},d)}{e(f_{i})}=r(f_{i},d)$, and that
$\sum_{i=1}^{I_{\ell}}\frac{p(f_{i},d)}{e(f_{i})}=r(d,-\ell),$
where $r(d,-\ell):=\#\mathcal{R}(d,-\ell)$.
9. Partial Zeta functions and Latimer’s work
In this Section, we will establish a link between the partial zeta functions $\zeta_{\xi_{i},\mathcal{P}}(s)$ of (14) and partial zeta functions attached to the quaternion algebra $B$. From Proposition 8.2 we have, for any $d\geq 1$, $n(\xi_{i},d)=r(f_{i},d)$, and from the previous discussion $r(f_{i},d)=p(f_{i},d)/e(f_{i})$. Hence,
$\zeta_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi_{i},N)N^{-s}=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}r(f_{i},N)N^{-s}=\\ =\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}p(f_{i},N)N^{-s}=\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P},\\ <u,v>=\mathcal{O}_{K}\end{subarray}}\frac{1}{f_{i}(u,v)^{s}},$
where the notation $(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}$ means that we exclude the pairs $(u,v)$ which give $\gcd(f_{i}(u,v),p)\neq 1$ for $p\in\mathcal{P}$. We now set
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s):=\zeta_{K,\mathcal{P}}(s)\zeta_{\xi_{i},\mathcal{P}}(s),$ (32)
where $\zeta_{K,\mathcal{P}}(s)$ is the Dedekind zeta function of $K$, with Euler factors above primes in $\mathcal{P}$ removed. We now have the following Lemma.
Lemma 9.1.
Assume that $K$ has class number one. Then
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}}\frac{1}{f_{i}(u,v)^{s}}.$
Proof.
This follows from factoring the possible common factors of $u$ and $v$ in the second series (compare also with [EGM97, Proposition 4.5]). ∎
To proceed further, we rely on the two papers of Latimer [Lat35, Lat36] on the relation between representation numbers of binary Hermitian forms and ideals in quaternion algebras.
Recall that in Section 7, we defined $B=K+\epsilon K$ and fixed the order $\mathfrak{o}=\mathcal{O}_{K}+\epsilon\mathcal{O}_{K}$ in $B$. For a left ideal $\mathfrak{U}$ in $\mathfrak{o}$, we may choose elements $\omega_{1},\omega_{2}\in\mathfrak{o}$ such that
$\mathfrak{U}=\mathcal{O}_{K}\omega_{1}+\mathcal{O}_{K}\omega_{2}.$
We call $\omega_{1},\omega_{2}$ a basis of $\mathfrak{U}$ with respect to $\mathcal{O}_{K}$ and write $\mathfrak{U}=[\omega_{1},\omega_{2}]$. If we write $\omega_{i}=g_{i1}+\epsilon\,\overline{g_{i2}}$, with $g_{ij}\in\mathcal{O}_{K}$, we define
$N(\mathfrak{U}):=\det\begin{pmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{pmatrix}$. (33)
(see [Lat35, Section 4]). In the case when $N(\mathfrak{U})$ is a positive integer, we call $\omega_{1},\omega_{2}$ a proper basis. Following [Lat35, Theorem 1], we will call $\mathfrak{U}$ a regular ideal if it possesses a proper basis. From [Lat35, Lemma 4], if $\mathfrak{U}$ is a regular ideal, every ideal in its (left) class is also regular. Hence, the term regular class is meaningful.
The main theorem in [Lat35, Theorem 3] reads as follows:
Theorem 9.2.
There is one-to-one correspondence between regular classes of left ideals in $\mathfrak{o}$ and $\textup{SL}_{2}(\mathcal{O}_{K})$-classes of positive definite binary hermitian forms.
We now write $\{\mathfrak{U}_{i}\}_{i=1}^{h_{1}}$ for representatives of regular classes of ideals in $\mathfrak{o}$ corresponding to the binary Hermitian forms $\{f_{i}\}_{i=1}^{h_{1}}$, where we recall that $f_{i}\in\Lambda[\ell,2^{-1}\mathbb{Z}]\subseteq\Lambda$. We first establish the following Proposition.
Proposition 9.3.
The ideals $\mathfrak{U}_{i}$, $i=1,\ldots,h_{1}$, are invertible.
Proof.
Let us consider a $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]\subseteq\Lambda$, and write $f_{i}=\begin{pmatrix}a&b\\ \bar{b}&c\end{pmatrix}$, with $a,c\in\mathbb{Z}$ and $b\in\mathcal{O}_{K}$. By the proof of [Lat35, Theorem 3], a representative of the class corresponding to this form is $\mathcal{L}:=[a,b+\epsilon]$. It suffices to show $\mathcal{L}$ invertible, as then every ideal in its class will be invertible (see [Voi21, Chapter 17]). Recall, $\epsilon^{2}=-k=b\bar{b}-ac<0$. Let $\mathcal{L}^{\prime}:=[a,\bar{b}-\epsilon]$, which is the ideal obtained by taking the standard involution of $B$. We compute
$\mathcal{L}\mathcal{L}^{\prime}=a^{2}\mathcal{O}_{K}+(a\bar{b}-a\epsilon)\mathcal{O}_{K}+(ab+\epsilon a)\mathcal{O}_{K}+(b\bar{b}-b\epsilon+\epsilon\bar{b}-\epsilon^{2})\mathcal{O}_{K}=\\ =a(a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K})+a\epsilon\mathcal{O}_{K},$
since $(b+\epsilon)(\bar{b}-\epsilon)=b\bar{b}-b\epsilon+\epsilon\bar{b}-\epsilon^{2}=b\bar{b}+k=ac$. Hence,
$\mathcal{L}\mathcal{L}^{\prime}=a(a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}+\epsilon\mathcal{O}_{K})$
We now claim that in order to establish that $\mathcal{L}$ is invertible with inverse $\frac{1}{a}\mathcal{L}^{\prime}$, it is enough to show that
$a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}=\mathcal{O}_{K}.$ (34)
Indeed, if this is true, we will have that $\mathcal{L}\left(\frac{1}{a}\mathcal{L}^{\prime}\right)=\mathfrak{o}$, so it would be enough to show that $\mathfrak{o}=\mathcal{O}_{L}(\mathcal{L})$. Following [Voi21, paragraph 16.5.18], we have
$\mathfrak{o}=\mathcal{L}\left(\frac{1}{a}\mathcal{L}^{\prime}\right)=\mathcal{O}_{L}(\mathcal{L})\mathcal{L}\left(\frac{1}{a}\mathcal{L}^{\prime}\right)=\mathcal{O}_{L}(\mathcal{L})\mathfrak{o}=\mathcal{O}_{L}(\mathcal{L}),$
because $\mathfrak{o}\subseteq\mathcal{O}_{L}(\mathcal{L})$, as $\mathcal{L}$ is an $\mathfrak{o}$-ideal. That is, $\mathfrak{o}=\mathcal{O}_{L}(\mathcal{L})$, which shows that $\mathcal{L}$ is invertible.
We now establish our claim in (34), i.e.,
$a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}=\mathcal{O}_{K}.$
We will make use of the assumption that $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$, and $\ell$ is an integer. Note that from [Shi06a, (2.18)], the first condition is equivalent to
$a\mathbb{Z}+c\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K})=\mathbb{Z},$
and the second implies that $b\in\mathcal{O}_{K}$ (see Lemma 8.1), which in turn gives
$a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}\subseteq\mathcal{O}_{K}.$
For the other direction, we write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}=\mathbb{Z}+\overline{\omega}\mathbb{Z}$, with $\omega$ as in (17). We then have
$\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}=(a\mathbb{Z}+c\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K}))+\omega(a\mathbb{Z}+c\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K}))=\\ =a\mathcal{O}_{K}+c\mathcal{O}_{K}+\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K})+\omega\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K})\subseteq a\mathcal{O}_{K}+c\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K},$
since $\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K})\subset b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}$ and similarly $\omega\textup{Tr}_{K/\mathbb{Q}}(b\mathcal{O}_{K})\subset b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}$. ∎
For a (left) ideal $I$ of $\mathfrak{o}$, we set $\mathcal{N}(I):=\#(\mathfrak{o}/I)$. We will now prove the following Lemma, which is given without proof as a footnote in [Lat35, page 441].
Lemma 9.4.
If $I$ is a regular left ideal in $\mathfrak{o}$, we have $N(I)^{2}=\mathcal{N}(I)$, where $N(I)$ is the norm of equation (33).
Proof.
First of all, if $\{a_{i}\},\{b_{i}\}$ are $\mathbb{Z}$-bases for $\mathfrak{o}$ and $I$ respectively, related by a base change matrix $C$, we have $\mathcal{N}(I)=|\det(C)|$. We write $\mathcal{O}_{K}=\mathbb{Z}+\mathbb{Z}\omega$ and then we have that a $\mathbb{Z}$-basis for $\mathfrak{o}$ is $\{1,\omega,\epsilon,\omega\epsilon\}$. Moreover, assume we take an $\mathcal{O}_{K}$-basis for $I$, given by $\theta_{1}=x+y\epsilon$ and $\theta_{2}=z+w\epsilon$, with $x,y,z,w\in\mathcal{O}_{K}$. A $\mathbb{Z}$-basis for $I$ is then given by $\{\theta_{1},\omega\theta_{1},\theta_{2},\omega\theta_{2}\}$. We write $x=a+b\omega$, $y=c+d\omega$, $z=e+f\omega$, and $w=g+h\omega$, with $a,b,c,d,e,f,g,h\in\mathbb{Z}$.
Consider first the case $K=\mathbb{Q}(\sqrt{-m})$ with $m\equiv 1,2\pmod{4}$. Then, using $\omega^{2}=-m$, $\overline{\omega}=-\omega$ and the fact that $\epsilon\omega=\overline{\omega}\epsilon=-\omega\epsilon$, we have that the matrix $C$ can be written as
$C=\begin{pmatrix}a&b&c&d\\ -mb&a&-md&c\\ e&f&g&h\\ -mf&e&-mh&g\end{pmatrix},\ m\equiv 1,2\pmod{4}$.
If $m\equiv 3\pmod{4}$ and using the fact that $\overline{\omega}=1-\omega$ and $\omega^{2}=\omega-(1+m)/4$, we get
$C=\begin{pmatrix}a&b&c&d\\ -(1+m)b/4&b+a&-d(1+m)/4&c+d\\ e&f&g&h\\ -(1+m)f/4&f+e&-h(1+m)/4&g+h\end{pmatrix},\ m\equiv 3\pmod{4}$.
In both cases, we have $N(I)=\det\begin{pmatrix}x&y\\ z&w\end{pmatrix}$. Using the fact that $I$ is a regular ideal and hence $N(I)$ is a positive integer, we can compute $N(I)^{2}=N(I)\overline{N(I)}=|\det(C)|=\mathcal{N}(I)$ for both cases, as claimed. ∎
Proposition 9.5.
Assume $K$ has class number $1$. For each $i=1,\cdots,h_{1}$, we have
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}I\in[\mathfrak{U}_{i}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$
where the summation is over all left ideals $I$ in $\mathfrak{o}$ in the class of $\mathfrak{U}_{i}$ that have norm not divisible by any prime in the set $\mathcal{P}$.
Proof.
For this, we rely on [Lat36, Theorem 1]. On the one hand, we have from Lemma 9.1 that
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}}\frac{1}{f_{i}(u,v)^{s}}=\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}d=1\\ (d,p)=1\ \forall p\in\mathcal{P}\end{subarray}}^{\infty}q(f_{i},d)d^{-s},$
where $q(f_{i},d):=\#\{(u,v)\in\mathcal{O}_{K}^{2}\backslash\mathcal{P}\mid f_{i}(u,v)=d\}$. From [Lat36, Corollary, p. 443] and because each $f_{i}$ is taken positive definite, we have that for each $d\geq 1$ with $\gcd(d,p)=1$ for all $p\in\mathcal{P}$
$q(f_{i},d)=e(f_{i})\cdot\#\{J\in[\mathfrak{U}_{i}]\textup{ with }N(J)=d\}.$
From this, the Proposition follows. ∎
We have stated the Proposition above in terms of left ideals in $\mathfrak{o}$, but the presence of the main involution of $B$ allows us to transfer the results also for right-ideals. Given a left ideal $I$ in $\mathfrak{o}$, we can obtain a right ideal $\bar{I}:=\{\bar{a}\,\,|\,\,a\in I\}$, where $\,\bar{}\,$ here denotes the main involution of $B$. Then one obtains a bijection between left and right ideals which preserves norms. Indeed, let us show that $N(\bar{I})=N(I)$. From [Voi21, Section 16.6.14], for any invertible ideal $J$, we have $\bar{J}J=N(J)\mathcal{O}_{R}(J)$, and $J\bar{J}=N(J)\mathcal{O}_{L}(J)$. By setting $J=\bar{I}$ in the first relation and $J=I$ in the second, we obtain
$I\bar{I}=\textup{nrd}(\bar{I})\mathcal{O}_{R}(\bar{I})\ \textup{ and }\ I\bar{I}=\textup{nrd}(I)\mathcal{O}_{L}(I).$
From [Voi21, Lemma 16.6.7], we have $\mathcal{O}_{R}(\bar{I})=\mathcal{O}_{L}(I)$, and hence we obtain that $\textup{nrd}(I)=\textup{nrd}(\bar{I})$. But $\textup{nrd}(I)=N(I)$, from our Lemma 9.4 and [Voi21, Theorem 16.1.3].
In particular, we may restate the above Proposition as follows.
Proposition 9.6.
Assume $K$ has class number $1$. For each $i=1,\cdots,h_{1}$, we have
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$
where the summation is over all right ideals $I$ in $\mathfrak{o}$ in the class of $\overline{\mathfrak{U}_{i}}$ that have norm not divisible by any prime in the set $\mathcal{P}$.
10. Main Theorem
In this Section, we are finally ready to state and prove our main Theorem. The first step is to identify everything in the quaternion algebra $B$. We have already done that for the zeta functions $\zeta_{\xi_{i},\mathcal{P}}(s)$ in the previous Section. As we have already mentioned in Section 7, $\Xi(A^{+}(W))=B$. Moreover, $\Xi(A^{+}(\Lambda)\cap A^{+}(W))=\mathfrak{o}$ from [Shi06b, (2.19b)]. Therefore, $\Xi(A^{+}(\Lambda\cap W))=\mathfrak{o}$ from [Mur14, Lemma 2.1].
By substituting (30) to (15), we obtain
$\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)\\ =L_{\mathcal{P}}(F;s-k+2)\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{P}}(\overline{f_{j}};s-k+5/2)^{-1}f_{j}(\tau(\zeta_{i}))=\\ =L_{\mathcal{P}}(F;s-k+2)\sum_{j=1}^{h_{2}}A_{f_{j}}L_{\mathcal{P}}(\overline{f_{j}};s-k+5/2)^{-1}\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3).$ (35)
Note that as has been observed in [Psy25, Theorem 9.3], we have $L_{\textup{Sug}}\left(\textbf{1};s-1/2\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L$-function we encounter in [Sug85] and [Shi99] respectively, as these are normalised differently. From now on, we use the normalisation of Shimura. Therefore, from now on, we will write $L_{\mathcal{P}}(f_{j},s-k+3)$ instead of $L_{\mathcal{P}}(f_{j},s-k+5/2)$. However, we still keep the normalisation of Sugano for $L_{\mathcal{P}}(F,s)$.
[…]
We set $S:=A^{+}(\Lambda\cap W)$ and $\mathcal{J}:=G^{+}(W)_{\mathbb{A}}\cap J$, and we select the finite set of primes $\mathcal{P}$ large enough so that $\mathcal{J}_{p}=(S_{p})^{\times}\cap G^{+}(W)_{p}$ for all $p\not\in\mathcal{P}$. We then consider the set $\mathcal{Y}:=\{y\in G^{+}(W)_{\mathbb{A}}\,\,|\,\,y_{p}\in\mathcal{Y}_{p}\textup{ for all }p\}$, where
$\mathcal{Y}_{p}=\begin{cases}G^{+}(W)_{p}\cap S_{p}\,\,\,\text{for}\,\,\,p\not\in\mathcal{P}\\ \mathcal{J}_{p}\,\,\text{otherwise}\end{cases}.$
We further define $\mathcal{Y}^{0}:=\{y\in\mathcal{Y}\,\,|\,\,\kappa(y)=\mathbb{Z}\}$, where $\kappa$ is defined in [Shi04, Section 22.8], and is called the capacity ideal. Moreover, we fix a set $\mathcal{A}\subset G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$ of pre-images for the elements $\{v_{l}\}_{l=1}^{h_{2}}$ under $\tau$. Assume $\{a_{l}\}_{l=1}^{h_{2}}$ are these elements. We then define
$\mathcal{Y}_{a_{\ell}}:=\{y\in\mathcal{Y}\,\,|\,\,\tau(y)\in H(\xi)_{\mathbb{Q}}\tau(a_{\ell})(H(\xi)_{\mathbb{A}}\cap C)\},$
and
$\mathcal{Y}_{a_{\ell}}^{0}:=\mathcal{Y}^{0}\cap\mathcal{Y}_{a_{\ell}}.$
Moreover, we set $Q:=\left\{q\in\prod_{p}\mathbb{Z}_{p}\,\,|\,\,q_{p}\neq 0,\,\,q-1\in p\mathbb{Z}_{p},\,\forall p\in\mathcal{P}\right\}$, and $R:=\mathcal{J}\cap\mathbb{Q}^{\times}_{\mathbf{h}}$, where $\mathbb{Q}^{\times}_{\mathbf{h}}$ denotes the finite adeles of $\mathbb{Q}$. We then have the following Lemma.
Lemma 10.1.
We have a bijection
$\mathcal{Y}_{a_{\ell}}/\mathcal{J}\longleftrightarrow Q/R\times\mathcal{Y}^{0}_{a_{\ell}}/\mathcal{J}.$
Moreover,
$\mathcal{Y}/\mathcal{J}=\bigsqcup_{\ell=1}^{h_{2}}\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ (36)
Proof.
First of all, from [Shi04, Lemma 22.9, (iii)], we have that there is a bijection between the set $\mathcal{Y}/\mathcal{J}$ and the set $Q/R\times\mathcal{Y}^{0}/\mathcal{J}$ by mapping a $y\in\mathcal{Y}$ to $(q,y_{0})$, where $y=qy_{0}$ with $q\in Q$ and $y_{0}\in\mathcal{Y}^{0}$. Moreover, from [Shi04, Lemma 22.10] we have that
$\mathcal{Y}^{0}/\mathcal{J}=\bigsqcup_{\ell=1}^{h_{2}}\mathcal{Y}^{0}_{a_{\ell}}/\mathcal{J}.$
The Lemma then follows by observing that $\tau(q)=1$ for all $q\in Q$, and hence $\tau(y)=\tau(y_{0})$ for $y=qy_{0}$ as before. ∎
We now state and prove the following Proposition.
Proposition 10.2.
Assume $K$ has class number $1$. We have the equality
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
where $\nu$ is the map of (3) and $\rho$ the map of Proposition 7.2.
Proof.
In order to establish the equality, we first use the fact that the map $\Xi$ of Proposition 5.4 identifies $A^{+}(W)$ with $B$. Using this identification, we have that the map $\nu$ is the reduced norm of the quaternion algebra $B$ (see, for example, [Shi04, Section 7.3]). In particular, we may view the right-hand side of the claimed equality as a partial zeta series of $B$. Under this identification, we have that $\mathcal{Y}_{p}=\mathfrak{o}_{p}\cap B_{p}^{\times}$ for all $p\not\in\mathcal{P}$. Moreover, we have $\mathcal{J}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p$ and hence in particular $\mathcal{Y}_{p}=\mathfrak{o}_{p}^{\times}$ for all $p\in\mathcal{P}$. Moreover the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ can be identified with locally principal right ideals in the order $\mathfrak{o}$ with the property that their image with respect to the map $\rho$ is the order corresponding to $w_{\ell}$ and locally are equal to $\mathfrak{o}_{p}$ for all $p\in\mathcal{P}$. Indeed, for an element $y=(y_{p})\in\mathcal{Y}$ we associate the locally principal ideal $I$ in $\mathfrak{o}$ such that $I_{p}=y_{p}\mathfrak{o}_{p}$ for all $p\not\in\mathcal{P}$, and $I_{p}=\mathfrak{o}_{p}$ for all $p\in\mathcal{P}$. By [Voi21, Main Theorem 16.7.7] such an ideal is invertible, and moreover every invertible (right) ideal in $\mathfrak{o}$ is locally principal. That is, the set $\mathcal{Y}_{a_{\ell}}/\mathcal{J}$ may be identified with the set of invertible right ideals in $\mathfrak{o}$ with the property that they are locally trivial for $p\in\mathcal{P}$, and their image with respect to the map $\rho$ is the order $w_{\ell}$. We now note that given an element $y=(y_{p})\in\mathcal{Y}$ corresponding to an ideal $I$, we have that $N(\nu(y)\mathbb{Z})=N(I)$. Indeed, by [Voi21, Main Theorem 16.7.7 (iv’)] and Lemma 9.4 above, we have that $N(I)=\textup{nrd}(I)$, where $\textup{nrd}(I)$ now equals $\textup{nrd}(I)=\textup{gcd}\{\nu(a)\,\,|\,\,a\in I\}$. Then, by [Voi21, Equation 16.3.4], we have that
$\textup{nrd}(I)=\prod_{p}N(\nu(y_{p})\mathbb{Z})=N(\nu(y)\mathbb{Z}),$
which concludes our claim. Finally, we also observe that the last equation gives that an invertible ideal $I$ of $\mathfrak{o}$ is locally trivial at a prime $p$ if and only if $\textup{gcd}(N(I),p)=1$. Putting all these together, we have that
$\sum_{\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}=\sum_{\rho(\overline{\mathfrak{U}_{i}})=w_{\ell}}\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$
where the first sum runs over a set of representatives of right ideals in $\mathfrak{o}$ which map to $w_{\ell}$ under the map $\rho$. Then Proposition 9.6 allows us to conclude the proof. ∎
We now define the space of automorphic forms on $G^{+}(W)_{\mathbb{A}}$ of weight $k$ with repsect to $\mathcal{J}$ as in [Shi04, (22.4a), (22.4b)]. We denote this space by $\mathcal{S}_{k}(\mathcal{J})$. Moreover, if $\psi$ is a Hecke character of $\mathbb{Q}$, we define
$S_{k}(\mathcal{J},\psi)=\left\{\mathbf{g}\in S_{k}(\mathcal{J})\mid\mathbf{g}(cx)=\psi(c)^{-1}\mathbf{g}(x),\ \forall c\in\mathbb{Q}_{\mathbb{A}}^{\times}\right\},$
as in [Shi04, (22.6)]. We also define the Hecke algebra $\mathfrak{R}(\mathcal{J},\mathcal{Y})$ in the sense of [Shi04, Section 8.16] and its action on $S_{k}(\mathcal{J,\psi})$ as follows. For an element $\tau\in\mathcal{Y}$, we take a finite subset $Y$ of $G^{+}(W)_{\mathbf{h}}$, so that
$\mathcal{J}\tau\mathcal{J}=\bigsqcup_{y\in Y}\mathcal{J}y.$ (37)
For $\mathbf{g}\in S_{k}(\mathcal{J},\psi)$ and $x\in G^{+}(W)_{\mathbb{A}}$, we then set
$(\mathbf{g}\mid\mathcal{J}\tau\mathcal{J})(x):=\sum_{y\in Y}\mathbf{g}(xy^{-1}),$ (38)
and extend the action linearly to $\mathfrak{R}(\mathcal{J},\mathcal{Y})$. Finally, for any $\mathbf{g}\in S_{k}(\mathcal{J},\psi)$, which is a Hecke eigenform, we define the Dirichlet series
$\mathfrak{T}(s,\mathbf{g}):=\sum_{\tau\in\mathcal{J}\backslash\mathcal{Y}/\mathcal{J}}N(\nu(\tau)\mathbb{{Z}})^{-s}\mu(\tau),$
where $\mu(\tau)\in\mathbb{C}$ is given by $\mathbf{g}\mid\mathcal{J}\tau\mathcal{J}=\mu(\tau)\mathbf{g}$.
We now give the following Lemma, which is similar to [Shi04, (22.13)], but here we consider left instead of right coset representatives.
Lemma 10.3.
Let $\mathbf{g}\in S_{0}(\mathcal{J},\mathbf{1})$, where $\mathbf{1}$ denotes the trivial character. We then have
$\mathfrak{T}(s,\mathbf{g})\mathbf{g}(x)=\sum_{y\in\mathcal{Y}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}\mathbf{g}(xy_{\mathbf{h}}),\,\,\forall x\in G^{+}(W)_{\mathbb{A}}.$
Proof.
First, as in [Shi04, p. 208], for each $\tau\in\mathcal{Y}$, we may find a set $Y_{\tau}$ such that
$\mathcal{J}\tau\mathcal{J}=\bigsqcup_{y\in Y_{\tau}}\mathcal{J}y=\bigsqcup_{y\in Y_{\tau}}y\mathcal{J},$
and so $\mathcal{J}\tau^{*}\mathcal{J}=\bigsqcup_{y\in Y_{\tau}}\mathcal{J}y^{*}=\bigsqcup_{y\in Y_{\tau}}\mathcal{J}\nu(y)y^{-1}$, where we have used the fact that $\mathcal{J}^{*}=\mathcal{J}$. Indeed, in our case the set $\mathcal{J}$ can be identified with the set $\prod_{p}\mathfrak{o}_{p}^{\times}$.
Using now the definition of the action of (38), we have
$(\mathbf{g}\mid\mathcal{J}\tau^{*}\mathcal{J})(x)=\sum_{y\in Y_{\tau}}\mathbf{g}(x\nu(y)^{-1}y)=\sum_{y\in Y_{\tau}}\mathbf{g}(xy),$
because $\mathbf{g}$ is of trivial character. Now we use the fact that $\mathcal{Y}^{*}=\mathcal{Y}$ since in our case $\mathcal{Y}_{p}=\mathfrak{o}_{p}$ for $p\not\in\mathcal{P}$, and $\mathfrak{o}_{p}^{\times}$ otherwise. But then we have that
$\mathfrak{T}(s,\mathbf{g})\mathbf{g}(x)=\sum_{\tau\in\mathcal{J}\backslash\mathcal{Y}/\mathcal{J}}N(\nu(\tau)\mathbb{{Z}})^{-s}\mu(\tau)\mathbf{g}(x)=\sum_{\tau\in\mathcal{J}\backslash\mathcal{Y}/\mathcal{J}}N(\nu(\tau^{*})\mathbb{{Z}})^{-s}\mu(\tau^{*})\mathbf{g}(x)=\\ =\sum_{\tau\in\mathcal{J}\backslash\mathcal{Y}/\mathcal{J}}N(\nu(\tau)\mathbb{{Z}})^{-s}(\mathbf{g}\mid\mathcal{J}\tau^{*}\mathcal{J})(x)=\sum_{y\in\mathcal{Y}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}\mathbf{g}(xy_{\mathbf{h}}),$
because $\mathcal{J}\tau\mathcal{J}=\bigsqcup_{y\in Y_{\tau}}y\mathcal{J}$. ∎
We now prove the following auxiliary Lemma.
Lemma 10.4.
For an eigenform $\mathbf{g}\in S_{0}(\mathcal{J},\mathbf{1})$, we have
$L(\mathbf{g},s)=L(\bar{\mathbf{g}},s).$
Proof.
First we note that the eigenvalues of $\mathbf{g}$ are real. This follows from the fact that the Hecke operators are self-adjoint with respect to the corresponding inner product (see, for example, [Shi04, (22.11a)], with $\psi$ trivial there). Hence, if we write $\mu(\tau)\in\mathbb{R}$ for the Hecke eigenvalue of $g$ with respect to the Hecke operator $\mathcal{J}\tau\mathcal{J}$, with $\tau\in\mathcal{Y}$, and consider a decomposition as in (37), we have
$\left(\overline{\mathbf{g}}|\mathcal{J}\tau\mathcal{J}\right)(x)=\sum_{y\in Y}\overline{\mathbf{g}(xy^{-1})}=\overline{\left(\mathbf{g}|\mathcal{J}\tau\mathcal{J}\right)(x)}=\overline{\mu(\tau)\mathbf{g}(x)}=\mu(\tau)\overline{\mathbf{g}}(x),$
That is, $\mathbf{g}$ and $\overline{\mathbf{g}}$ have the same system of eigenvalues. But the $L$-function is attached to the system of eigenvalues, and hence they are the same. ∎
We are now ready to prove the following Theorem.
Theorem 10.5.
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. For each $j=1,\cdots,h_{2}$, we have
$\zeta_{K,\mathcal{P}}(s-k+3)\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3)=f_{j}(1)L_{\mathcal{P}}\left(f_{j};s-k+3\right).$ (39)
Proof.
[…]
$\zeta_{K,\mathcal{P}}(s-k+3)\sum_{i=1}^{h_{1}}f_{j}(\tau(\zeta_{i}))\zeta_{\xi_{i},\mathcal{P}}(s-k+3)=\sum_{l=1}^{h_{2}}f_{j}(v_{l})\sum_{\begin{subarray}{c}i\textup{ such that }\\ \tau(\zeta_{i})=v_{l}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s-k+3).$
[…]
$\sum_{\ell=1}^{h_{2}}f_{j}(\tau(a_{\ell}))\left(\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s-k+3)\right)=f_{j}(1)L_{\mathcal{P}}(g_{j};s-k+3).$ (40)
We now use Lemma 10.3 with $\mathbf{g}=g_{j}=f_{j}\circ\tau$ and set $x=1\in G^{+}(W)_{\mathbb{A}}$. We then have
$\mathfrak{T}(s,g_{j})g_{j}(1)=\sum_{y\in\mathcal{Y}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}g_{j}(y_{\mathbf{h}}).$
Now, because $\tau$ is a homomorphism, we have $g_{j}(1)=f_{j}(\tau(1))=f_{j}(1)$. From (36) in Lemma 10.1, we can rewrite the above as
$\mathfrak{T}(s,g_{j})f_{j}(1)=\sum_{\ell=1}^{h_{2}}\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s}g_{j}(y_{\mathbf{h}})=\sum_{\ell=1}^{h_{2}}f_{j}(\tau(a_{\ell}))\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
because if $y\in\mathcal{Y}_{a_{l}}$, then $\tau(y)\in H(\xi)\tau(a_{l})(H(\xi)_{\mathbb{A}}\cap C)$ and hence $f_{j}(\tau(y_{\mathbf{h}}))=f_{j}(\tau(a_{l}))$.
It is now true from [Shi04, (22.12b)] that $\mathfrak{T}(s,g_{j})=L(g_{j},s)$. Proposition 10.2 now gives that
$\sum_{\begin{subarray}{c}i\textup{ such that }\\ \rho(u_{i})=w_{\ell}\end{subarray}}\widehat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{y\in\mathcal{Y}_{a_{\ell}}/\mathcal{J}}N(\nu(y)\mathbb{Z})^{-s},$
and hence (40) follows. ∎
We can finally state and prove the Main Theorem of this paper.
Theorem 10.6.
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. Assume $F\in S_{k}(\Gamma_{S})$ is a Hecke eigenform, and consider the Poincaré series $\mathcal{P}_{k,D,r}$ of (10) with $D=-\ell q$ and $r\in L$ such that $(-\ell q,r)\in\widetilde{\textup{supp}}(L,\phi)$ of (9). Then, if $\mathcal{P}$ is a finite set of primes, including the primes described in [Psy25, Theorem 8.2], we have
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=A(\xi)L_{\mathcal{P}}(F;s-k+2),$
Proof.
By substituting (39) from Theorem 10.5 in (35), using also Lemma 10.4, we obtain
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=C\times L_{\mathcal{P}}(F;s-k+2),$
[…]
Remark 10.1.
The fact that the constant $C$ equals $A(\xi)$ can also be deduced from the fact that $\{f_{j}\}_{j=1}^{h_{2}}$ is an orthonormal basis for $V(\xi)$.
We now write $\mathcal{V}_{k}$ for the span of $\mathcal{P}_{k,D,r}$ for all valid choices of $D$ and $r$, as in the Theorem above. We then obtain the following Corollary.
Corollary 10.1.
Let $G\in\mathcal{V}_{k}$. Then, for a finite set of primes $\mathcal{P}$ sufficiently large we have that
$\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,G}(s)=\langle\phi_{1},\psi_{1}\rangle L_{\mathcal{P}}(F;s-k+2),$
where $\phi_{1}$ and $\psi_{1}$ are the first Fourier-Jacobi coefficients of $F$ and $G$, respectively.
Proof.
From the Theorem above, it follows that $\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)D_{F,G}(s)$ is equal to $L_{\mathcal{P}}(F;s-k+2)$ up to a constant, when we take $\mathcal{P}$ large enough to apply our theorem for each Poincaré series in the linear combination that makes up $G$. This constant must then be $\langle\phi_{1},\psi_{1}\rangle$, by comparing the coefficients of $1^{-s}$ in the expansion of the Dirichlet series. ∎